### v1

# Monotone and Separable Set Functions: Characterizations and Neural Models

Abstract
Motivated by applications for set containment problems, we consider the following fundamental problem: can we design set-to-vector functions so that the natural partial order on sets is preserved, namely $S\subseteq T\text{ if and only if }F(S)\leq F(T)$. We call functions satisfying this property Monotone and Separating (MAS) set functions. We establish lower and upper bounds for the vector dimension necessary to obtain MAS functions, as a function of the cardinality of the multisets and the underlying ground set. In the important case of an infinite ground set, we show that MAS functions do not exist, but provide a model called MASNet which provably enjoys a relaxed MAS property we name “weakly MAS” and is stable in the sense of Holder continuity. We also show that MAS functions can be used to construct universal models that are monotone by construction and can approximate all monotone set functions. Experimentally, we consider a variety of set containment tasks. The experiments show the benefit of using our MASNet model, in comparison with standard set models which do not incorporate set containment as an inductive bias. Our code is available in GitHub.
1 Introduction
A multiset $\{\!\!\{x_{1},\ldots,x_{n}\}\!\!\}$ is an unordered collection of vectors, where order does not matter (like sets) and repetitions are allowed (unlike sets). In recent years, there has been increased interest in neural networks that can map multisets to vectors, with applications for physical simulations [1], processing point clouds [2], and graph neural networks [3]. Another important application of multiset-to-vector maps is set-containment search [4, 5, 6]. Here the goal is to check whether a given multiset $S$ is (approximately) a subset of $T$, and this is often carried out by learning a multiset-to-vector mapping $F$, and then checking whether $F(S)\leq F(T)$ (element-wise), in which case, one deduces that $S$ is a subset of $T$. In this paper, we look into this problem from a theoretical perspective.
Monotone and Separable functions:
We begin our analysis with some definitions: we say that a function $F$ mapping sets to vectors is monotone if $S\subseteq T\subseteq V$ implies that $F(S)\leq F(T)$, and we will say that $F$ is separable if $F(S)\leq F(T)$ implies that $S\subseteq T$. When $F$ is simultaneously monotone and separable, we call it a MAS function. When $F$ is a MAS function, we can safely test whether $F(S)\leq F(T)$, and this will be fully equivalent to checking whether $S\subseteq T$.
Motivation for MAS functions:
It is natural to ask why we’re interested in functions that are MAS instead of just being monotone or just separable. For this, we begin with a couple of real-life applications based on the set-containment task: (I) In recommendation system design, one problem is to recommend item having a particular set of features, $S$. Here, we represent each item as a set of corresponding features $T$, and the problem is to find all sets $T$ such that approximately, $S\subseteq T$. (II) Text entailment where we are given a small query sentence $q$ nd the goal is to find the set of corpus items $c$ from a large corpus $C$, where $q\implies c$ or $c$ entails $q$. Here, $q$ and $c$ are typically represented as sets of contextual embeddings ($S$ for $q$ and $T$ for $c$) and the entailment problem can be cast as the problem of checking if $S\subset T$.
In a neural network setting, such problems require us to design a function $F$ such that the "order" between $F(S)$ and $F(T)$ can serve as a boolean test for whether $S$ is a subset of $T$. For this, $F$ needs to satisfy two conditions: (A) $F(S)\leq F(T)$ implies $S\subseteq T$ (B) $F(S)\not\leq F(T)$ implies $S\not\subseteq T$. Here, we like to emphasize that, for high accuracy, $F$ needs to satisfy both conditions A and B as above. Monotone functions satisfy condition (B) without necessarily satisfying condition (A). As a result, if we predict $S\subseteq T$ based on $F(S)\leq F(T)$ when $F$ is monotone but not separable, it will give large number of false positives. Similarly, if $F$ is separable but not monotone, then $F$ satisfies condition (A) without satisfying condition (B). Thus, using such for checking $S\subseteq T$ results in a large number of false negatives. Hence, both monotonicity and separability are necessary to develop an accurate test for set-containment
Related works:
Monotone set functions have been extensively studied from theoretical perspective, e.g., in the context of theory of capacities [7], game theory [8], fuzzy systems [9], combinatorial auctions [10, 11, 12] and learning theory [13, 14, 15, 16, 17, 18]. However, to the best of the knowledge, the notion of functions which are jointly monotone and separable multiset (MAS), is entirely new, albeit some related works on partial order of sets can be traced in mathematics literature, such as [19]. Our goal in this paper is to understand whether it is possible to construct MAS multiset functions, and if so, develop differentiable models that are MAS by construction and are appropriate for multiset containment tasks.
Our inquiry is also motivated by the study of multiset injective functions. This has been extensively studied in recent works [20, 3], including characterization of the vector dimension needed to enable injectivity [21, 22, 23] and the development of differentiable injective multiset models [24, 25, 26]. In this work, we aim to attain similar results for the more challenging problem of MAS multiset functions (more details in Subsection 2.1).
An additional goal of this work is stability: in most learning scenarios, we are looking for $S$, which is only approximately a subset of $T$. Accordingly, we would like to design functions $F$ that are not only MAS but also stable, in the sense that when $S$ is approximately a subset of $T$, then $F(S)$ is approximately dominated by $F(T)$.
Summary of our goal
In summary, in this work, we aim to characterize (MAS) set functions that output finite-dimensional set representations and subsequently design neural networks for such functions. At a high level, we seek to address the following questions: (1) What are the conditions for existence of a finite dimensional MAS functions? (2) What are the permissible relaxations to monotonicity and separability if the existential conditions identified in (1) are not satisfied? (3) What are the possible neural architectures for these functions? (4) Are these functions stable? We expect that the design of trainable models with built-in guarantees of monotonicity, separability, and stability will significantly enhance the inductive bias for set containment applications.
1.1 Main Results
We address the goal specified in the previous subsection by providing a detailed characterization and neural architecture for monotone and separable set functions. Our main results are as follows:
Existential characterization of MAS functions
Our first objective is to determine whether it is possible to obtain a MAS set function with finite output dimension $m$. We begin our analysis in the case where set elements are assumed to be taken from a finite ground set $V$, and show that in this case a MAS function exists if, and only if, $m\geq|V|$. Next, we add the assumption that the cardinality of the input multisets is bounded by $k$, and provide lower and upper bounds for $m$ in this case. When the ground set is infinite, we show that MAS set functions do not exist.
Weakly MAS functions
In most applications, the ground set is $V=\mathbb{R}^{D}$, and our analysis shows that in this case, MAS functions do not exist in general. To address this, we introduce the notion of weakly MAS functions by extending the set functions to parametric set functions $F(S,w)$, where $w$ is a parameter. Weakly MAS functions requires that (i) $F(:,w)$ is monotone, for all $w$, and (ii) for any $S\not\subset T$, there exists at least one $w\in\mathcal{W}$ such that $F(:,w)$ separates $S$ and $T$. We explain how, by choosing suitable activations, it is possible to construct simple deep sets [20] models which are weakly MAS, and name the resulting model MASNet.
Stability
We define the notion of stability as discussed above, separability notions are Boolean and fail to capture graded distance. We would like to guarantee that $F(T)$ is ’almost’ larger than $F(S)$ when $S$ is ’almost’ a subset of $T$. To address this, we propose a novel asymmetric set distance and present a Holder separability condition [27] which ensures stability in terms of this asymmetric distance.
Monotone Functions
Finally, we show that MAS functions can be used to provide universal models for computing set-to-vector functions. Thus, MAS functions may be a useful concept also outside of the set containment setting.
Experiments
We provide a number of experiments showing that for set containment problems, using our weakly MAS model leads to stronger results in comparison with standard multiset models like DeepSets and SetTransformer, which do not incorporate MAS considerations as an inductive bias.
Summary of contributions
Our main contributions in this paper are: (1) We introduce the novel notion of monotone and separating (MAS) multiset functions. (2) We discuss lower and upper bounds for the embedding dimension for MAS functions, and suggest a model MASNet, which is weakly MAS. (3) We show the stability of MASNet. (4) Experimentally, we prove the effectiveness of MASNet for set containment tasks.
2 Existential characterization of MAS functions
We begin by stating the notation we will use for our paper.
Notation
We denote by $V$ to be the ground set and $\mathcal{P}_{<\infty}(V)$ to be the collections of all multisets from $V$ with finite cardinality $S=\{\!\!\{x_{1},\cdots,x_{s}\}\!\!\}$ where $x_{i}\in V$. We denote the space of all multisets with at most $k$ elements by $\mathcal{P}_{\leq k}(V)$. We use $[d]$ to denote the set $\left\{1,\cdots,d\right\}$. Given two vectors $v,u\in\mathbb{R}^{d}$, we say $v\leq u$ if $v[i]\leq u[i]$ for all $i\in[d]$. Also, for an element $v\in V$ and a multiset $S\in\mathcal{P}_{<\infty}(V)$, we use $c_{S}(v)$ to denote the number of times $v$ occurs in $S$. We say $S\subseteq T$ if $c_{S}(v)\leq c_{T}(v)$ for all $v\in V$.
Proofs
All proofs for the results stated in the paper are provided in Appendix 8.
In this section, our goal is to understand when an MAS function $F:\mathcal{P}_{\leq k}(V)\to\mathbb{R}^{m}$ or $F:\mathcal{P}_{<\infty}(V)\to\mathbb{R}^{m}$ exists, and if it does, what is the smallest $m$ for which such a mapping exists. We denote this minimal dimension as $m^{\star}(V,k)$, and $m^{\star}(V,\infty)$, respectively.
2.1 Prologue: Relation to injectivity
Our inquiry is related to the notion of injective multiset functions, as every MAS function $F$ is, in particular, injective. Indeed, suppose for two sets $S,T$ we have $F(S)=F(T)$. Then, we have both $F(S)\leq F(T)$ and $F(T)\leq F(S)$. Hence, we have both $S\subseteq T$ and $T\subseteq S$ by separability, which implies $S=T$. Thus, injectivity follows from separability.
Injective multiset functions $f:\mathcal{P}_{\leq k}(V)\to\mathbb{R}^{m}$ are attainable even for $m=1$, providing that $V$ is finite or even infinite [20, 3, 21, 24]. As we will see, attaining the strong condition of MAS multiset functions requires high dimension, and in some cases (infinite $V$), it does not even exist. This can be seen even in the following very simple example:
Example 1.
Assume the ground set consists of only two elements $V=\{0,1\}$, and we only look for multisets of cardinality $\leq k=1$. In this case, there are only three multisets in the space $\mathcal{P}_{\leq 1}(V)$: the empty set, $S=\{0\}$, and $T=\{1\}$. As discussed previously, there exists a multiset function $F:\mathcal{P}_{\leq 1}(V)\to\mathbb{R}$ which is injective. However, no such function can be separated. This is because neither $S$ nor $T$ is a subset of the other. However, because real numbers are totally ordered, we have either $F(S)\leq F(T)$ or $F(T)\leq F(S)$, violating the separability condition.
We note that the monotonicity can be satisfied bu using functions of the form $F(S)=\sum_{x\in S}f(x)$ for non-negative $f$. Thus, achieving separation is harder than achieving monotonicity.
The key reason for the non-existence of a scalar MAS set function is that multisets (or sets) are not totally ordered, but scalars are. However, this issue does not arise for multidimensional set functions, where vectors, similar to sets, are partially ordered. So, when mapping to vectors, when and how can MAS functions be constructed? We will now discuss this.
2.2 Existence of MAS functions: finite ground set, unbounded cardinality
We study MAS functions $F:\mathcal{P}_{<\infty}(V)\to\mathbb{R}^{m}$, where input multisets may have an arbitrarily large finite cardinality, and the ground set is finite $|V|=n$. In this case, we can show that the smallest possible dimension $m^{\star}(V,k)$ of a MAS function is exactly $n$:
Theorem 1 (Dimension of MAS Function).
For a finite ground set $V$ of size $n$, there exists a MAS function $F:\mathcal{P}_{<\infty}(V)\to\mathbb{R}^{n}$. In addition, any MAS function must have a dimension of at least $n$. In other words, $m^{\star}(V,\infty)=n$.
Proof Sketch
The construction of a monotone embedding with $n=|V|$ simply uses one-hot encoding. Namely, we identify $V$ with $[n]$. For every $S\in\mathcal{P}_{<\infty}(V)$, we define
$F(S)=\sum_{s\in S}e_{s}\in\mathbb{R}^{n}\textstyle$ (1)
where $e_{s}\in\mathbb{R}^{n}$ is the vector with $e_{s}[s]=1$ and $e_{s}[j]=0$ for all $j\neq s$. It’s clear $F$ satisfies all conditions. For the second part, assume we have an embedding of dimension $m$, for every output dimension $i\in[m]$, there is a "maximal singleton element" $v^{*}_{i}\in V$ such that $F(\left\{v^{*}_{i}\right\})[i]\geq F(\left\{v\right\})[i],\forall v\in V$. The value of the set function $F$ applied to $T=\left\{v^{*}_{1},\ldots,v_{m}^{*}\right\}$ will dominate any singleton, due to the monotonicity of $F$. However, if $m<|V|$, then we can select a $u\in V\setminus T$ and then $F(\left\{u\right\})\leq F(T)$, which contradicts separability.
2.3 Existence of MAS functions: finite ground set, finite cardinality
We now consider the case where the ground set $V$ is finite, as before, but now the multisets have bounded cardinality $k<|V|$. Indeed, in many practical applications, input multiset cardinality is much smaller than $|V|$, even when $|V|$ is large. In this setting, we show that we can get a lower embedding dimension $m$ than in Theorem 1. In fact, we show the minimal output dimension $m$ of a MAS function $F:\mathcal{P}_{\leq k}(V)\to\mathbb{R}^{m}$ can scale logarithmically with the size of the ground set $|V|$. However, this comes at the price of an exponential dependence on $k$:
Theorem 2 (Upper Bound on $m^{\star}(V,k)$).
Let $k<n$ be natural numbers, and let $V$ be a ground set with $|V|=n$. Then there exists a MAS function $F:\mathcal{P}_{\leq k}(V)\to\mathbb{R}^{m}$ with embedding dimension $m=(k+2)^{k+2}\log(n)$.
Proof sketch
We construct the MAS function by taking the one-hot embedding $F:\mathcal{P}_{\leq k}(V)\to\mathbb{R}^{n}$ from (1), and then applying $m$ random projections defined by vectors in $\mathbb{R}^{n}$ with non-negative entries. The non-negativity ensures monotonicity, and we prove that for the stated value of $m$, the probability of achieving a MAS function using this procedure is strictly positive.
The theorem shows that $m^{\star}(V,k)\leq(k+2)^{k+2}\log(n)$. We now give two lower bounds on the embedding dimension: we show that $m$ must depend at least linearly on $k$, and at least double logarithmically on $n$:
Theorem 3 (Lower bounds on $m^{\star}(V,k)$).
Let $k\geq 2$ and $n$ be natural numbers, and let $V$ be a ground set with $|V|=n$. Then the smallest possible dimension $m^{\star}(V,k)$ of a MAS function satisfied $m^{\star}(V,k)\geq\log_{2}(\log_{3}n)$. Moreover, if $k\leq\frac{n-1}{2}$ then $m^{\star}(V,k)\geq 2k$
Proof Sketch:
To obtain the first lower bound, we consider the sequence of singleton-set embeddings $\mathcal{M}:=(F(\left\{1\right\}),\ldots,F(\left\{n\right\})$. By the Erdös-Szekers theorem [28], every scalar sequence with $n$ elements has a monotone subsequence of length $\sim n^{\frac{1}{2}}$. Applying this recursively to vectors of length $m$ gets us a monotone subsequence in all $m$ coordinates of the vector, of length $\sim n^{\frac{1}{2^{m}}}$. If this subsequence is of length $\geq 3$ then we will have three distinct elements $v_{1},v_{2},v_{3}\in V$ such that $F(\{v_{1}\})[i]\leq F(\{v_{2}\})[i]\leq F(\{v_{3}\})[i]\text{ or }F(\{v_{3}\})[i]\leq F(\{v_{2}\})[i]\leq F(\{v_{1}\})[i],\forall i\in[m]$ Monotonicity implies that $F(\{v_{2}\})[i]$ is dominated by $F(\{v_{1},v_{3}\})[i]$, for all $i\in[m]$, thus contradicting separability. It follows that MAS existence can only happen when $n^{\frac{1}{2^{m}}}\leq 2$, which leads to the double logarithmic lower bound.
For the lower bound in terms of $k$, the argument is similar to the proof of  Theorem˜1, where we select a "maximal singleton element" $v^{*}_{i}$ for every dimension $i\in[m]$. Since m<|V|1m<|V|-1, we can find elements $\left\{u_{1},u_{2}\right\}$ disjoint from the collection $\bigcup_{i=1}^{m}\left\{v^{*}_{i}\right\}$, and WLOG $F(\left\{u_{1}\right\})[i]\geq F(\left\{u_{2}\right\})[i]$ for at least half of the indices in $[m]$. We then construct $S=\left\{u_{2}\right\}$ and $T$ to be union of $u_{1}$ and all the $v^{*}_{i}$ for all dimensions $i$, where $F(\left\{u_{1}\right\})[i]\leq F(\left\{u_{2}\right\})[i]$. Then monotonicity implies $F(S)\leq F(T)$, but $S\not\subseteq T$. This implies $k\leq|T|-1\leq m/2$, as otherwise separability of $F$ in $\mathcal{P}_{\leq k}(V)$ is violated.
2.4 Non-existence of MAS functions: for infinite ground set
For most practical applications, we deal with infinite (often uncountable) ground sets, for example $V=\mathbb{R}^{d}$. Thus, we would like to analyze whether MAS functions exists when $|V|=\infty$. The answer to this question is negative:
Corollary 4.
Given a ground state $V$ with $|V|=\infty$, and $k\geq 2$, there does not exist a MAS function $F:\mathcal{P}_{\leq k}(V)\to\mathbb{R}^{m}$ for any finite $m\in\mathbb{N}$.
This corollary is a simple consequence of the result in Theorem 3 that the embedding dimension cannot be larger than $\log\log|V|$ when $V$ is finite. We note that this result makes the minimal assumption $k\geq 2$. In Appendix 8 we show this assumption is necessary, and that in the degenerate case $k=1$ there is a MAS function $F:\mathcal{P}_{\leq 1}(V)\rightarrow\mathbb{R}^{2}$ for V=[1,1]V=[-1,1].
Summary
In this section, we provided lower and upper bounds for the smallest value $m^{\star}(V,k)$ for which an MAS function $F:\mathcal{P}_{\leq k}(V)\to\mathbb{R}^{m}$ exists. Our results are summarized in Table 1. We note that some of these results require weak assumptions, which are stated in the theorems but not in the table.
Ground set size $\qquad\qquad$ $\infty |V|=n=|V|=n$$\infty$
Input multiset size $\quad$ $\infty k=k$$\infty$ $k=1$ $k\geq 2$
$m^{\star}(V,k)\leq\min\{n,(k+2)^{k+2}\log(n)\}$, $m^{\star}(V,k)\geq\max\{2k,\log_{2}\log_{3}n\}$ $m^{\star}(V,k)=n$ $m^{\star}(V,k)=2$ Not possible
Table 1: This table summarizes the lower and upper bounds for the smallest value $m^{\star}(V,k)$ for which an MAS function $F:\mathcal{P}_{\leq k}(V)\to\mathbb{R}^{m}$ exists, where $k$ is maximal set cardinality and $n$ denotes the cardinality of the ground set $V$.
3 Relaxations of MAS functions
Our definitions so far did not address differential parameterized set functions $F(S,w)$, which are the natural object of interest for set learning applications. One natural way to address this could be to require that there exists $w$ such that $F(\bullet,w)$ is a MAS function; however, since our theoretical results show that MAS functions do not exist when $|V|=\infty$, this requirement is too restrictive. Instead, we retain monotonicity as a hard constraint for every parameter $w\in\mathcal{W}$, while treating separability as an existential condition over the parameter space.
3.1 The notion of weakly MAS functions
Formally, given a ground set $V$, and a probability space $(\mathcal{W},\mathcal{B},\mu)$, we consider parametric set functions $F:\mathcal{P}_{<\infty}(V)\times\mathcal{W}\to\mathbb{R}^{m}$, where $F(S,\bullet)$ is measurable for all $S$. We then define weakly MAS functions as:
Definition 5 (weakly MAS function).
The set function $F:\mathcal{P}_{<\infty}(V)\times\mathcal{W}\to\mathbb{R}$ is a weakly MAS function if the following two conditions are satisfied:
(1) Pointwise monotonicity: For any $w\in W$ and $S\subseteq T$, we have that $F(S,w)\leq F(T,w)$.
(2) Weak separability: If $S\not\subseteq T$, then there exists $w\in\mathcal{W}$ such that $F(S;w)>F(T;w)$.
Which condition, monotonicity or separability, should be relaxed?
In our definition, we relax the separability requirement so that different multiset pairs can be separated by different parameters, while we maintain strict requirements of pointwise monotonicity. This is because, as discussed in Example 1, constructing monotone set functions is straightforward, even in one dimension. In contrast, separability is a much stronger and more difficult condition to satisfy. This is a key reason why MAS functions do not exist when $|V|=\infty$. Indeed, as shown in Appendix 8, there can be no continuous separable set function even without the monotonicity assumption when the ground set is $V=\mathbb{R}^{D}$. Therefore, separability is the natural condition for relaxation. Moreover, monotonicity remains a useful and often necessary constraint in many applications.
Are scalar deep sets weakly MAS?
Equipped with the new notion of weakly MAS functions, we ask the question of how to construct such a class of functions. As a first step to achieve this goal, we look into a simple instance of DeepSets [20]. DeepSets defines a set function by applying an ’inner’ MLP $\mathrm{M}_{1}$ to each set element, summing over the result, and then applying an ’outer’ MLP $\mathrm{M}_{2}$. We consider the case where $\mathrm{M}_{1}$ is a shallow MLP $\mathrm{M}_{1}(x)=\sigma(Ax+b)$ which gives us models of the form
$F(S;(A,b))=\textstyle\mathrm{M}_{2}\big(\sum_{x\in S}\sigma(Ax+b)\big)$ (2)
To ensure monotonicity for every parameter choice $A,b$, we will require that $\mathrm{M}_{2}$ is a monotone vector-to-vector function, and the activation $\sigma$ is non-negative. While these choices automatically ensure monotonicity, they may not lead to weak MAS functions as for many activations there will be no separation:
Proposition 6.
If the deep set model in (2) is implemented with a vector-to-vector monotonously increasing function $\mathrm{M}_{2}$ and a non-negative activation $\sigma$, then $F(\bullet;(A,b))$ is monotone for every $(A,b)$. If in addition $\sigma$ is monotone (increasing or decreasing), then there exists $S\not\subseteq T$ with $F(S;(A,b))\leq F(T;(A,b))$ for all $A,b$.
Proof idea The monotonicity is rather straightforward. To prove lack of separation when $\sigma$ is monotone, choose $x\neq y,\ z=\frac{1}{2}(x+y),\ S=\{z\},\ T=\{x,y\},$ and then $Az+b$ is the average of $Ax+b$ and $Ay+b$, which can be used to show that $F\left(T;(A,b)\right)\geq F\left(S;(A,b)\right)$ for all $A,b$.
Are Set Transformers weakly MAS?
Now we ask the quedtion if Set Transformers [29] are weakly MAS. We notice in the following that Set Transformers are not even monotone in the following result. For this, we consider Set Transformer with sum-pooling, namely $F(S;W_{Q},W_{K},W_{V})=\text{SumPool}(\text{Attn}(S))$. Explicity, this is defined for a give multiset $X=\left\{\mathbf{x}_{1},\cdots,\mathbf{x}_{n}\right\}$ by first defining query, key, values per-point: $\mathbf{q}_{i}=W_{Q}\mathbf{x}_{i},\mathbf{k}_{i}=W_{K}\mathbf{x}_{i},\mathbf{v}_{i}=W_{V}\mathbf{x}_{i}$, and then taking a weighted average, defined by weights: $\alpha_{i,j}=\frac{e^{\mathbf{q}_{i}\cdot\mathbf{k}_{j}}}{\sum_{s}e^{\mathbf{q}_{i}\cdot\mathbf{k}_{s}}}$ to obtain a permutation-invariant function $F(S;W_{Q},W_{K},W_{V}):=\sum_{i,j}\alpha_{i,j}\mathbf{v}_{i}$
Proposition 7.
Let $V=\mathbb{R}^{d}$ and let $F:\mathcal{P}_{<\infty}(V)\to\mathbb{R}^{D}$ be a Set Transformer defined by $F(S;W_{Q},W_{K},W_{V}):=\text{SumPool}(\text{Attn}(S))$. Then $F$ is not a point-wise monotone function.
Proof Idea:
We consider full rank matrices $W_{Q}=W_{K}$ and any non-zero $W_{V}$. We choose $\mathbf{x}_{1}\in\mathbb{R}^{d}$ such that $\mathbf{v}_{1}:=W_{V}\mathbf{x}_{1}\neq 0$, and consider the sets $S=\left\{\mathbf{x}_{1}\right\},T=\left\{\mathbf{x}_{1},\mathbf{0}_{d}\right\}$. Then, $F(S;W_{Q},W_{K},W_{V})=\mathbf{v}_{1}$ and $F(T;W_{Q},W_{K},W_{V})=(\alpha_{1,1}+\alpha_{1,2})\mathbf{v}_{1}$. Note that, $\alpha_{1,1}+\alpha_{1,2}>1$ and thus, for any negative $\mathbf{v}_{1}[j]$, we have $F(T;W_{Q},W_{K},W_{V})[j]<F(S;W_{Q},W_{K},W_{V})[j]$ and thus, monotonicity is violated.
Thus, unlike Deep Sets, it’s not obvious how to make Set Transformers monotone, let alone weakly separable. Thus, in the following sections, we shall focus on obtaining weakly-MAS functions from DeepSet-like models only.
3.2 The Hat activation class
Note that most commonly used non-negative activation functions, such as $\mathrm{ReLU}$ or sigmoid, are monotonically non-decreasing and therefore, from Proposition 6, using them as $\sigma$ above will fail weak separability. Here, we consider a novel class of activation functions we call ’Hat Activations’, which will all make $F$ in Eq. (2) weakly MAS.
Definition 8 (The Hat activation function).
We call a function $\sigma:\mathbb{R}\to\mathbb{R}$ a hat activation if it is (a) non-negative (b) compactly supported (c) not identically zero and (d) continuous.
Examples of hat functions are the third and fourth functions in Figure 2.
Figure 2: Using a multiset model as in (2) with activations which are not always non-negative, like $\sigma=\mathrm{Tanh}$ will not be monotone. ReLU will be monotone, but to be weakly MAS, two layers are required. TRI and more general hat functions are weakly MAS even with a single layer.
When $\sigma$ is a hat function, Equation˜2 is weakly MAS, even when the output of $F$ is taken to be scalar. For this result, we will need to require that $\mathrm{M}_{2}$ is strictly monotone. This can be handled by simply setting $\mathrm{M}_{2}$ to be the identity.
Proposition 9.
If $\mathrm{M}_{2}$ is strictly monotone increasing, and $\sigma$ is a hat activation function, then $\mathrm{M}_{2}(\sum_{x\in S}\sigma(a^{\top}x+b))$ is weakly MAS function.
Proof idea
If $S$ is not a subset of $T$, there is an element $s$ whose multiplicity in $S$ is larger than its multiplicity in $T$. We can then choose $a,b$ so that $s$ is in the support of $\sigma(ax+b)$ but the elements of $T$ are not.
3.3 Slightly larger ReLU networks
So far, we have considered simple models as in (2) which apply a single linear layer before applying activation and summation. In this case, we saw that $\mathrm{ReLU}$ and other activations do not create weakly MAS functions, and defined the notion of hat activations, which are weakly MAS functions. But what happens if we use $\mathrm{ReLU}$ activations but allow deeper networks?
We can use our previous analysis to show that very small $\mathrm{ReLU}$ networks with two layers can already be MAS functions. To see this, we consider a candidate from the hat function class, namely the $\mathrm{TRI}$ function from Figure 2, and observe that: $\mathrm{TRI}(x)=\mathrm{ReLU}\left(\mathrm{ReLU}(2x)+\mathrm{ReLU}(2x-2)-\mathrm{ReLU}(4x-2)\right)$ We use this to deduce the following result:
Proposition 10.
Given $V\subseteq\mathbb{R}^{d}$, we consider affine transformations $\mathcal{A}_{2}:\mathbb{R}^{d}\to\mathbb{R}^{3},\mathcal{A}_{1}:\mathbb{R}^{3}\to\mathbb{R}$. Thus, here $\mathcal{A}_{2}(t):=A_{2}t+b_{2}$ and $\mathcal{A}_{1}(z)=a_{1}^{\top}z+b_{1}$, where $A_{2}\in\mathbb{R}^{3\times d},b_{2}\in\mathbb{R}^{3},a_{1}\in\mathbb{R}^{3},b_{1}\in\mathbb{R}$ are the respective parameters. Then, the set functions of the form
$F:\mathcal{P}_{<\infty}(V)\to\mathbb{R},\textstyle F(S;\mathcal{A}_{1},\mathcal{A}_{2})=\mathrm{M}_{2}\left(\sum_{x\in S}\mathrm{ReLU}\circ\mathcal{A}_{1}\circ\mathrm{ReLU}\circ\mathcal{A}_{2}(x)\right)$ (3)
are weakly MAS functions.
Our results so far are summarized in Figure 2. If $\sigma$ is not a non-negative function, then $F$ will not be monotone. We can attain weakly MAS functions with a two-layer ReLU network, or with a one-layer network with hat activations.
3.4 Asymmetric distance induced Holder separability
Why weak separability is not sufficient
While weak separability is a key relaxation for weakly MAS functions (Definition 5), it has two key limitations: (1) Separation should ideally hold over a non-negligible subset of $\mathcal{W}$, not just a single point $w$; (2) it treats separability as a boolean condition, ignoring approximate containment. We would like to guarantee that if $S$ is almost a subset of $T$, then $F(S,w)$ will be close to dominated by $F(T,w)$. This motivates us to create an asymmetric (pseudo) metric to quantify the size of set difference and create stronger notions of separability based on that.
Asymmetric set distance
For a ground set $V\subseteq\mathbb{R}^{d}$ and two sets $S,T\in\mathcal{P}_{<\infty}(V)$ such that $|S|\leq|T|$. We define the asymmetric distance from $S$ to $T$ as the Earth Mover Distance (EMD) between $S$ and the subset of $T$ that is closest to $S$ with the same cardinality as $S$.
$\Delta(S,T)=\min_{T^{\prime}\subseteq T:|S|=|T^{\prime}|}\mathrm{EMD}(S,T^{\prime})$ (4)
$\Delta(\bullet,\bullet)$ captures graded set containment: if $S\not\subseteq T$, but is "very close" to a subset of $T$ in EMD metric, then $\Delta(S,T)$ is small. Moreover, $\Delta(S,T)=0\iff S\subseteq T$. Based on $\Delta$, we propose a newer separability notion for parametric functions.
Lower Hölder separability of set functions
Davidson and Dym [27] introduced Hölder continuity through expectation over parameters. Unlike their symmetric EMD, we define Hölder separability using the asymmetric distance $\Delta$(4).
Definition 11 (Lower Hölder separability).
Let $V\subseteq\mathbb{R}^{d}$ be the ground set and $(\mathcal{W},\mathcal{B},\mu)$ be a probability space, and a constant $\lambda>0$. A parametric set function $F(\bullet,w):\mathcal{P}_{\leq k}(V)\to\mathbb{R}^{m}$ with $w\sim\mu(\cdot)$ is $\lambda$ lower Hölder separable if there exists $c>0$ such that:
$\mathbb{E}_{w\sim\mu(\cdot)}\lVert\left[F(S;w)-F(T;w)\right]_{+}\rVert_{1}\geq c\cdot\Delta(S,T)^{\lambda},\\ \ \text{for all }S,T\in\mathcal{P}_{\leq k}(V)$ (5)
3.5 Hölder separable Set functions
Hölder separability using Hat activation
Proposition 9 shows how to construct weakly MAS functions using Hat activations. We now show that these constructions are also lower Hölder, under additional weak assumptions on the hat functions and $M_{2}$:
Theorem 12 ($F$ is lower Hölder).
[…]
Probability of successful separation
Given $S,T\subseteq V$, we can consider $F(S;(a,b,c))-F(T;(a,b,c))]_{+}$ to be a real-valued non-negative random variable, where the randomness is over the parameter space $(a,b,c)\in\mathcal{W}=\mathbb{R}^{d}\times\mathbb{R}\times\mathbb{R}$ equipped with a probability measure. The above result on lower Hölder stability gives us a lower bound on the probability of a parameter tuple $(a,b,c)$ separating two non-subsets $S,T$ proportional to the set distance $\Delta(S,T)$. Moreover, the functions we have considered so far are scalar-valued set functions. By considering independent copies of the parameters across multiple dimensions, we can increase the probability of separation, since $F(S),F(T)$ are not separated iff $F(S)[i]\leq F(T)[i]$ across each embedding dimension $i$. We formalize the above observations as follows:
Theorem 13 (Probability bounds on separation).
Let $V\subseteq\mathbb{R}^{d}$ and $\sigma$ be as in Theorem 12, and let $A\in\mathbb{R}^{m\times d},b\in\mathbb{R}^{m},c\in\mathbb{R}^{m}$ whose $m$ columns (respectively entries) are drawn independently from the distribution on $a_{j},b_{j},c_{j}$ described in Theorem 12, and consider the function
$F(S;A,b,c)=\sum_{x\in S}\sigma\left(c^{-1}\odot(Ax+b)\right).$ (6)
Then there exists $C>0$, so that for all $S\not\subseteq T$, $\mathbb{P}\left(F(S)\leq F(T)\right)\leq\left(1-C\Delta^{2}(S,T)\right)^{m}.$
In Equation (6) $c^{-1}$ stands for the elementwise inverse and $\odot$ stands for elementwise multiplication.
Note that the probability of failed separation goes to zero exponentially as $m$ increases, and also becomes smaller as $\Delta(S,T)$ increases. This supports the idea that a larger measure of parameters separates sets with larger asymmetric distance and that separability becomes easier with larger embedding dimensions (which we saw previously in the existential results)
Hölder separable set functions using ReLU
As seen in Proposition 10, one can construct weakly MAS functions with a two layer neural network with $\mathrm{ReLU}$ activation. Our proof used that two layer $\mathrm{ReLU}$ networks can basically "simulate" a Hat function. Under the light of the above results on Hölder separability, one can argue that two-layer $\mathrm{ReLU}$ networks would also have such guarantees. Also, we show in Appendix 8 that under certain structure on the ground set $V$ (for example, $V$ being a hypersphere), even with one layer $\mathrm{ReLU}$ networks, functions of the form in  Equation˜2 will also have Lower Hölder separability guarantees.
Upper Lipschitz bounds
In Appendix. 8 we complement our results on lower Holder stability by showing that the construction in Theorem 13 is also upper Lipschitz in an appropriate sense.
3.6 Monotone Universality and the role of $M_{2}$
Besides set containment-based applications, there are other interesting scenarios where one would like to construct monotone multiset-to-vector functions. The following theorem shows that on finite ground sets, the combination of a MAS function and a universal monotone vector-to-vector (such as [30]) can approximate all monotone multiset-to-vector functions.
Theorem 14 (Universality).
Let $V$ be a finite ground state, and let $F:\mathcal{P}_{\leq k}(V)\to\mathbb{R}^{m}$ be a MAS function. Then for every multiset-to-vector monotone function $f:\mathcal{P}_{\leq k}(V)\to\mathbb{R}^{s}$, there exists a vector-to-vector monotone function $M:\mathbb{R}^{m}\to\mathbb{R}^{s}$ such that $F(S)=M\circ f(S)$.
4 MASNet: Neural Modeling of MAS functions
Now, our goal is to leverage our theoretical analysis to design neural network-based multiset-to-vector models that preserve monotonicity and weak separability. Motivated by the formulation in DeepSets and our analysis so far, we wish to design neural set functions of the form:
$\textsc{MASNet}(S)=\textstyle M_{\theta_{2}}\left(\sum_{x\in S}\sigma\left(M_{\theta_{1}}(x)\right)\right)$ (7)
In our analysis, we required $M_{\theta_{2}}$ to be a monotone vector-to-vector function. In most of our experiments, we enforce this simply by choosing $M_{2}(x)=x$, but this can also be enforced by using monotone activations and non-negative parameters.
MASNet-Hat
To enforce weakly MAS in Equation˜7,our results from Prop. 9 and Theorem 12 suggests using a Hat activation as $\sigma$. In our experiment, we do it by not choosing $\sigma$ to be a specific hat activation, but rather the parametric form:
$\sigma_{\alpha,\beta,\gamma}(x)=\mathrm{ReLU}\left(\frac{x-\alpha}{\gamma\cdot\beta}\right)+\mathrm{ReLU}\left(\frac{x-(\alpha+\beta)}{(1-\gamma)\cdot\beta}\right)-\mathrm{ReLU}\left(\frac{x-(\alpha+\gamma\cdot\beta)}{\gamma\cdot(1-\gamma)\cdot\beta}\right)$ (8)
For all $\alpha$, $\beta>0$, and $\gamma\in(0,1)$, $\sigma_{\alpha,\beta,\gamma}$ is a hat function (Definition 8) with support $\alpha,\alpha+\beta]$ and peak at $\alpha+m\beta$. Examples include the third and fourth functions in Figure 2. Applying this to an $m$-dimensional output with independent $\alpha,\beta,\gamma$ per dimension yields $3m$ parameters total.
MASNet with $\mathrm{ReLU}$
Motivated by our theoretical results from  Proposition˜10 of 2-Layer $\mathrm{ReLU}$ networks being weakly MAS and from our discussion of $\mathrm{ReLU}$ networks being Hölder separable (under appropriate assumptions) in Section 3.5, we propose use $\mathrm{ReLU}$ in MASNet, which case $M_{\theta_{1}}$ will have to be a network with at least 2 layers to ensure weakly MAS as in Proposition˜10. This gives one more way to design MASNet and we refer to this option as MASNet-ReLU.
Other variants of MASNet
Definition 8 establishes a general hat function class extending beyond the piecewise linear functions as in (8). We present MASNet-INT in Appendix 9, which achieves universal approximation of the Hat class by modeling its derivative via neural networks and approximating the integral. Additionally, using the specific hat activation $\mathrm{TRI}$ (Figure 2) as $\sigma$ yields an alternative formulation called MASNet-$\mathrm{TRI}$.
Which MASNet to choose and when?
We have given multiple recipes of MASNet in the above discussion, and the question arises: which one to choose? We show via our experiments: For parameter-constrained scenarios requiring single-layer $M_{\theta_{1}}$ networks, MASNet-Hat provides optimal separability (both theoretically and empirically). For deeper $M_{\theta_{1}}$ ($\geq 2$ layers), both MASNet-ReLU and Hat-based variants perform similarly, with MASNet-ReLU often providing more stable training and better performance overall.
5 Experiments
We evaluate the MASNet variants (Section 4) on synthetic, text and point-cloud datasets to characterize monotonicity and separability. Specifically, we focus on (exact and approximate) set containment. Given sets $S$ and $T$, with binary label $y(S,T)\in\{0,1\}$ indicating (exact or approximate) set containment, we evaluate how accurately MASNet predicts $y(S,T)$.
5.1 Set Containment
We perform experiments on a synthetically generated dataset, four text datasets, and one image dataset. In each case, we split the dataset into 5:2:2 train, test, and dev folds. We minimize the following fixed-margin hinge loss that enforces vector dominance, to train the parameters of MASNet.
$\sum_{S,T}(1-y(S,T))\min_{i\in[m]}\left[F(S)[i]-F(T)[i]+\delta\right]_{+}+y(S,T)\max_{i\in[m]}\left[F(S)[i]-F(T)[i]+\delta\right]_{+}\vskip-5.69054pt$ (9)
$|S|$ $|T|$ DS ST M-ReLU M-Hat 1 2 0.98 0.98 0.99 0.99 1 10 0.59 0.59 0.99 0.99 10 30 0.89 0.80 1.00 1.00 10 100 0.54 0.57 0.98 0.99
Table 3: Accuracy on synthetic data
Analysis on synthetic datasets
To generate our synthetic dataset, we first sample a target set $T\subset\mathbb{R}^{d}$, where each element $x\in T$ is drawn from $\mathcal{N}(0,\bm{I})$, i.i.d. Given $T$, we compute $S$ with a fixed size $|S|=s$ as follows. We first obtain $S$ with positive labels $y(S,T)=1$ by drawing a subset from $T$, uniformly at random, without replacement. To generate $S$ with $y(S,T)=0$, we sample $s$ from $\mathcal{N}(0,\bm{I})$ independently. We chose $d=4$ and the set embeddings have $m=256$ dimensions, the class-ratio of subsets to non-subsets was taken to be 1:1.
We compare two variants of MASNetMASNet-ReLU (M-ReLU) and MASNet-Hat (M-Hat)— against Deep Sets (DS) [20] and Set Transformer (ST) [29]. Table 3 summarizes the results
Figure 4: Acc vs $|T|$ for 1-layer MLP
for varying values of $|S|$ and $|T|$. We observe that: (1) Both variants of our method outperform the baselines; (2) the baselines degrade significantly as $|T|$ increases (with fixed $|S|$), as predicting separability becomes harder as the gap between $|T|$ and $|S|$ increases; (3) The performance of M-ReLU and M-Hat is comparable. We now compare 1-layer pointwise models ending in $\mathrm{ReLU}$ and Hat activation respectively, followed by aggregation: we call them 1layer-$\mathrm{ReLU}$ and 1layerM-Hat. We see in Fig. 4 that: (1) 1layerM-Hat performs significantly than 1layer-$\mathrm{ReLU}$, validating the results from Propn. 6. (2) acc. declines for both as $|T|$ increases, consistent with the intuition that larger $|T\setminus S|$ gaps make the task harder.
Set containment on Text datasets
We evaluate on MSWEB, MSNBC, and Amazon Registry datasets, which exhibit natural set containment from user behavior. Each datapoint is a wordbag; we compute BERT embeddings for all unique words to form the ground set. $(S,T)$ pairs are sampled from these (Appendix 10), and labeled via set containment. For inexact containment, small
Model Bedding Feeding MSWEB MSNBC DeepSets 0.51 0.50 0.88 0.67 SetTransformer 0.77 0.79 0.90 0.94 FlexSubNet 0.88 0.85 0.92 0.91 Neural SFE 0.52 0.53 0.88 0.66 MASNet-ReLU 0.98 0.98 0.99 0.97 MASNet-Hat 0.94 0.93 0.97 0.95
Table 5: Acc in text datasets with 1:1 ratio
Gaussian noise is added to $S$ without changing $y(S,T)$. We now compare the variants of MASNet: MASNet-ReLU and MASNet-Hat against several baselines, DeepSets [20], Set Transformers [29], FlexSubNet [31], Neural SFE [32], for inexact set containment on the datasets: MSWEB, MSNBC, Amazon-Feeding, Amazon-Bedding. Gaussian noise with $0.01$ std was used to generate noisy pairs. The class ratio of subsets to non-subsets was taken to be 1:1. Here, set embeddings have dim $m=50$, and set elements have dim $d=768$. The results in table 5 show: (1) MASNet-ReLU and MASNet-Hat have similar acc., with MASNet-ReLU being marginally better (2) Both are generally better than baselines.
Set containment on point clouds:
We use ModelNet40 [33], a dataset of 3D CAD models across
$|S|\rightarrow$ $128$ $256$ $512$ DeepSets 0.52 0.51 0.50 SetTransformer 0.62 0.51 0.50 FlexSubNet 0.84 0.75 0.60 Neural SFE 0.52 0.51 0.50 MASNet-ReLU 0.98 0.94 0.72 MASNet-Hat 0.87 0.81 0.65
Table 6: Performance on Point cloud for different values of $|S|$ with 1:1 class ratio.
40 categories. For each object in category $C_{1}$, we sample 1024 points to form $T$. For true subsets, we sub-sample $S\subseteq T$ and for non-subsets, $S$ is sampled from a different category $C_{2}$. Small white noise(std=0.005) is added to $S$ for the task of inexact containment. We use a pointcloud encoder(such as pointnet) followed by a set-to-vector embedding model to produce $m=50$ dimensional embeddings, and each point in the dataset is a 3D coordinate with $d=3$ dimensions. The class ratio of subsets to non-subsets was taken to be 1:1. Results are in table 6. We observe that: (1) MASNet-ReLU and MASNet-Hat are generally better than baselines, with MASNet-ReLU being slightly better. (2) The accuracy of MASNet decreases as $|S|$ decreases, with the separability becoming hard for monotone models.
5.2 Monotone Set function approximation
Model MAE
DeepSets 0.01024 $\pm$ 0.00030
SetTransformer 0.01026 $\pm$ 0.00033
MASNet-ReLU 0.01023 $\pm$ 0.00107
MASNet-Hat 0.00959 $\pm$ 0.00016
Table 7: Monotone function approximation
In this experiment, we examine if a scalar variant of MASNet can approximate a monotone function $F^{*}$, by getting trained from set, value pairs in the form of $\left\{(S,F^{*}(S))\right\}$. Once trained, we measure $|MAS(S)-F^{*}(S)|$ for each $S$ and average over them to compute MAE. We compare MASNet against baselines: DeepSets, SetTransformer. We trained using MSE loss, and report test MAE results in Table 7 where we see: (1) MASNet-Hat is best performing, followed by MASNet-ReLU (2) Both are better than baselines.
6 Conclusion and Future Work
In this work, we study the design of set functions that are useful in set containment through Monotone and Separating (MAS) properties. We derive bounds on the embedding dimensions required for MAS functions and show their nonexistence in infinite domains. To address this, we introduce a relaxed model, MASNet, which satisfies a weak MAS property and is provably stable. Experiments demonstrate that MASNet outperforms standard models on set containment tasks by leveraging monotonicity as an inductive bias. It would be interesting to check if our method can be used for applications in graphs, e.g., subgraph matching. Another potential direction is to consider other possible relaxations of separability.
7 Limitations
Our work provides a theoretical analysis of monotonicity and separability for set functions in the context of set containment, and introduces neural models that outperform standard set-based architectures such as DeepSets and Set Transformers on set containment tasks. Our work has the following limitations, addressing which are interesting directions for future work:
  • After proving the impossibility of constructing MAS functions over infinite ground sets, we proposed parametric set functions defined with respect to a probability measure over the parameter space. We provided probabilistic guarantees: for non-subsets $S,T$, if their asymmetric distance $\Delta(S,T)$ is large, then a randomly sampled function from the parameter space will separate them with high probability. However, these guarantees hold only in the randomized setting—analogous to guarantees for neural networks at random initialization—and do not extend to models after training via gradient-based optimization.
  • Our universality result, which shows that MAS functions composed with monotone vector-to-vector mappings can approximate all monotone set-to-vector functions, was only established for finite ground sets. Extending this result to infinite ground sets remains an open problem.
  • We currently do not generalize our analysis to the subgraph isomorphism problem, which is a natural and strictly harder extension of the set-containment task.
Acknowledgment
Nadav and Yonatan were supported by ISF grant 272/23, Soutrik and Abir were supported by Amazon and Google Research Grant; as well as Bhide Family Chair Endowment Fund.
APPENDIX
8 Proofs of the technical results
8.1 Proofs of existential results on MAS functions
See 1
Proof.
The construction of a monotone embedding with $n=|V|$ simply uses one-hot encoding. Namely, we identify $V$ with $[n]$. For every $S\in\mathcal{P}_{<\infty}(V)$, we define
$F(S)=\sum_{s\in S}e_{s}\in\mathbb{R}^{n}$ (10)
where $e_{s}\in\mathbb{R}^{n}$ is the vector with $e_{s}[s]=1$ and $e_{s}[j]=0$ for all $j\neq s$. We now show it satisfies all conditions: Be $S\subseteq T$, then by definition, $\forall v\in V,c_{S}(v)\leq c_{T}(v)$, then $F(S)=\sum_{s\in S}e_{s}\leq\sum_{t\in T}e_{t}=F(T)$ thus $F(S)\leq F(T)$. On the other hand, if $F(S)\leq F(T)$ then $\forall v\in V,c_{S}(v)\leq c_{T}(v)$, and thus $S\subseteq T$.
For the second direction, assume there exists MAS $F:V\to\mathbb{R}^{m}$ for $m\leq n-1$. For every output dimension $i\in[m]$, there is a "maximal singleton element" $v^{*}_{i}\in V$ such that $F(\left\{v^{*}_{i}\right\})[i]\geq F(\left\{v\right\})[i],\forall v\in V$. The value of the set function over the union of such $v^{*}_{i}$-s across $i\in[m]$ will dominate any singleton that is disjoint from this collection, thanks to the monotonicity of $F$. As $m<|V|$, then $T:=\bigcup_{i=1}^{m}\left\{v^{*}_{i}\right\}$ does not cover $V$. Thus, we can select an element disjoint from this collection named $s\in V$, which gives disjoint sets $S=\{s\},T$ with $F(S)\leq F(T)$, which contradicts separability. ∎
See 2
Proof.
[…]
$h_{j}(S)\geq h_{j}(\tilde{S})>h_{j}(\tilde{T})\geq h_{j}(T).$
Given an extreme pair $(S,T)$ we consider the set of all ’bad’ $a$, namely
$B_{S,T}:=\{a\in[0,1]^{n},\langle f(S),a\rangle>\langle f(T),a\rangle\}$
Our first goal will be to bound the Lebesgue measure of this set. Recall that $S$ is a singleton $S=\{s\}$, and $T$ does not contain this singleton. Accordingly,
$B_{S,T}=\{a\in[0,1]^{n}|\quad a_{s}\leq\sum_{t\in T}a_{t}\}$
We will compute the probability of the complement of this set, namely vectors $a$ such that $a_{s}>\sum_{t\in T}a_{t}$. First denote by
$E_{S,T}=\{a\in[0,1]^{n}|\forall t\in T,a_{t}<\frac{a_{s}}{k}\}$
Note, that
$E_{S,T}\subseteq\{a\in[0,1]^{n}|\quad a_{s}>\sum_{t\in T}a_{t}\}$
So let’s compute the measure of $E_{S,T}$:
$\mathrm{Leb}(E_{S,T})=\mathrm{Leb}(\{a\in[0,1]^{n}|\forall t\in T,a_{t}\leq\frac{a_{s}}{k}\})=\int_{0}^{1}(\frac{a_{s}}{k})^{k}da_{s}=\frac{1}{k^{k}\cdot(k+1)}\geq\frac{1}{(k+1)^{k+1}}$
So,
$\mathrm{Leb}(B_{S,T})\leq 1-\mathrm{Leb}(E_{S,T})\leq 1-\frac{1}{(k+1)^{k+1}}$
Now, consider the set
$B^{m}_{S,T}=\{a_{1},\ldots,a_{m}\in B_{S,T}\},$
Then we have
$\mathrm{Leb}(B^{m}_{S,T})\leq\left(1-\frac{1}{(k+1)^{k+1}}\right)^{m}$
The vectors $a_{1},\ldots,a_{m}$ will not define a monotone embedding, if there exists an extreme pair $S,T$ which is in $B^{m}_{S,T}$. The probability of this happening can be bounded by a union bound. There are $n\cdot{n-1\choose k}$ extreme pairs, and for simplicity we will replace this number with a larger but simpler expression $n^{k+2}>n\cdot{n-1\choose k}$. The union bound will then give us
$\mathrm{Leb}\{a_{1},\ldots,a_{m}\text{ do not define a monotone embedding }\}\leq n^{k+2}\left(1-\frac{1}{(k+1)^{k+1}}\right)^{m}$
It is sufficient to show that this expression is smaller than $1$, for our $m$, which, by taking a logarithm, is equivalent to requiring that
$(k+2)\ln\left(n\right)+m\ln\left(1-\frac{1}{(k+1)^{k+1}}\right)<0$
Using Taylor’s expansion it can be shown that $\ln(1-x)<-x$ for $x\in(0,1)$, and as a result it is sufficient to choose $m$ so that
$(k+2)\ln\left(n\right)-\frac{m}{(k+1)^{k+1}}<0$
or equivalently
$m>(k+1)^{k+1}\cdot(k+2)\cdot\ln(n).$
For convenience we slightly enlarge this lower bound to obtain $m\geq(k+2)^{k+2}\ln(n)$
See 3
Proof.
The proof is partially inspired by the proof in [19].
Lower bound on $k$:
We want to show the lower bound of $m^{\star}(V,k)\geq\min\left(2k,n-2\right)$. As in the proof of Theorem˜1, we select a "maximal singleton element" $v^{*}_{i}\in V$ for every dimension $i\in\left\{1,2,\cdots,m^{\star}(V,k)\right\}$. Thus we must have:
$\text{for each }i\in[m^{\star}(V,k)]:\ F(\left\{v^{*}_{i}\right\})[i]\geq F(\left\{v\right\})[i],\forall v\in V$
Now, if $m^{\star}(V,k)\leq|V|-2$, we can find two elements $\left\{u_{1},u_{2}\right\}$ disjoint from the collection $\mathcal{V}:=\bigcup_{i=1}^{m^{\star}(V,k)}\left\{v^{*}_{i}\right\}$. Without loss of generality, we can assume that $F(\left\{u_{1}\right\})[i]\geq F(\left\{u_{2}\right\})[i]$ for most indices in $m^{\star}(V,k)]$. We can then construct a set $T$ containing $u_{1}$ and all the $v^{*}_{i}$ for all indices $i$ for which $F(\left\{u_{1}\right\})[i]<F(\left\{u_{2}\right\})[i]$, thus we define:
$T:=\left\{u_{1}\right\}\cup\left\{v_{j}^{*}:F(\left\{u_{1}\right\})[j]<F(\left\{u_{2}\right\})[j]\right\}$
We then have that:
$F(\left\{u_{2}\right\})[i]\leq\max\{F(\left\{u_{1}\right\})[i],F(\left\{v^{*}_{i}\right\})[i]\}\leq F(T)[i],\forall i\in[m^{\star}(V,k)]$
where the last inequality uses the monotonicity of $F$. This implies that $F(S)\leq F(T)$ However, $S$ is not a subset of $T$. It follows that $k\leq|T|-1\leq m^{\star}(V,k)/2$, as otherwise separability of $F$ in $\mathcal{P}_{\leq k}(V)$ is violated. Hence, we get that: $m^{\star}(V,k)\geq 2k$
Lower bound on $n$:
For this proof, we will use the Erdös-Szekeres theorem, which in particular states, for a natural $N\geq 2$, that a sequence of real numbers with $\geq N^{2}$ elements has a subsequence of cardinality $N$ which is monotone (either monotonously increasing or monotonously decreasing). Assume we have $F:\mathcal{P}_{\leq k}(V)\rightarrow\mathbb{R}^{m}$ a monotone embedding, where $V=[n]$. Assume by way of contradiction that:
$\log_{2}(\log_{3}n)>m\text{ or equivalently }n>3^{2^{m}}$
Consider the first coordinate and look at the sequence
$F(\{1\})[1],\ldots,F(\{n\})[1]$
This is a real-valued sequence, and so there is a monotone subsequence $a_{1}<a_{2}<\ldots,<a_{\ell}$ of the original sequence $1,\ldots,n$, with $\ell=\sqrt{3^{2^{m}}}=3^{2^{m-1}}$, such that:
[…]
Next, we consider the second coordinate of this subsequence, namely $F(a_{1})[2],\ldots,F(a_{\ell})[2]$ and obtain a new subsequence $\{a_{j}\}$ such that both the sequences: $\{F(a_{j})[1]\}$ and $\{F(a_{j})[2]\}$ are ordered monotonically, and the size of the subsequence is the square root of the previous one, namely $3^{2^{m-2}}$. After doing this $m$ times, we have a subsequence of size $3$. Namely, we have three distinct elements $u,v,w\in V$ such that for all $i=1,\ldots,m$,
$\text{either }F(\{u\})[i]\leq F(\{v\})[i]\leq F(\{w\})[i]\text{ or }F(\{u\})[i]\geq F(\{v\})[i]\geq F(\{w\})[i]$
It follows that for all $i$, we have:
$F(\{v\})[i]\leq\max\left(F(\{u\})[i],F(\{w\})[i]\right)\leq F(\{u,w\})[i]$
Thus, $F(\{v\})\leq F(\{u,w\})$, which is a contradiction since $\{v\}$ is not contained in $\{u,w\}$. ∎
A refined lower bound:
Now, we may extend the above proof idea to generalize the lower bound on $m^{\star}(V,k)$, which is useful specially when $|V|>>k$. In such cases, the following gives a tighter lower bound on $m^{\star}(V,k)$ which is quadratic in $k$.
Lemma 15.
For all $\ell\in[k]$, we have $m^{\star}(V,k)\geq\min\left(|V|-\ell,\ell k+2\ell-\ell^{2}\right)$. More specifically, if $\ell=\frac{k+2}{2}$, then $m^{\star}(V,k)\geq\min\left\{|V|-\frac{k+2}{2},\left(\frac{k+2}{2}\right)^{2}\right\}$
Proof.
Consider any $\ell\in[k]$. Like in the proof just before, we consider a "maximal singleton element" $v^{*}_{i}\in V$ for every dimension $i\in\left\{1,2,\cdots,m^{\star}(V,k)\right\}$ and obtain the collection $\mathcal{V}:=\bigcup_{i=1}^{m^{\star}(V,k)}\left\{v^{*}_{i}\right\}$. Let us assume $m^{\star}(V,k)\leq|V|$-$\ell$. Thus, we can produce the collection $\mathcal{U}:=\left\{u_{1},u_{2},\cdots,u_{\ell}\right\}$ disjoint from $\mathcal{V}$, i.e $\mathcal{V}\cap\mathcal{U}=\emptyset$. For each $u_{q}\in\mathcal{U}$, let $m_{q}$ be the number of co-ordinates $j\in[m^{\star}(V,k)]$ such that $F(\left\{u_{q}\right\})[j]>F(\left\{u_{r}\right\})[j],\forall r\in[\ell]\setminus\left\{q\right\}$. Clearly, $\sum_{q\in[\ell]}m_{q}\leq m^{\star}(V,k)$, thus, $\exists\ell_{0}\in[\ell]\setminus$ such that $m_{\ell_{0}}\leq\frac{m^{\star}(V,k)}{\ell}$. Like before, we consider the sets $S=\left\{u_{\ell_{0}}\right\}$. We construct $T$ as follows: we select $\mathcal{U}\setminus\left\{u_{\ell_{0}}\right\}$ and take union with all $\left\{v^{*}_{j}\right\}$ for each dimension $j$ in which $F(\left\{u_{\ell_{0}}\right\})[j]>F(\left\{u_{r}\right\}),\forall r\in[\ell]\setminus\left\{\ell_{0}\right\}$. Since $\left\{v^{*}_{j}\right\}$ is the maximal singleton for the dimension $j$, it follows that $F(\left\{u_{\ell_{0}}\right\})[j]\leq F(\left\{v^{*}_{j}\right\})[j]$. Hence, if we define:
$T=\left(\mathcal{U}\setminus\left\{u_{\ell_{0}}\right\}\right)\bigcup\left\{v^{*}_{j}:F(\left\{u_{\ell_{0}}\right\})[j]>F(\left\{u_{r}\right\}),\forall r\in[\ell]\setminus\left\{\ell_{0}\right\}\right\}$
[…]
Now, we already have a sufficient upper bound on $m^{\star}(V,k)$ from Theorem˜2 that’s $O\left(k^{k+2}\log(n)\right)$. Now, when $n>>k$ we have that $O\left(k^{k+2}\log(n)\right)<n-O(k)$, in which case we may apply $\ell=\frac{k+2}{2}$ in the above lemma. And since we already know that $O\left(k^{k+2}\log(n)\right)<n-O(k)$, we thus get that $m^{\star}(V,k)<n-O(k)$. Thus, in this case the above lemma gives us $m^{\star}(V,k)\geq\left(\frac{k+2}{2}\right)^{2}$, which is a tigher quadratic lower bound in $k$. ∎
When $k=1$
In the degenerate case where only a single set element is allowed, $k=1$, it is possible to construct a MAS function even for the uncountable ground set V=[1,1]V=[-1,1] by defining $F:\mathcal{P}_{\leq 1}([-1,1])\rightarrow\mathbb{R}^{2}$ via
$F(S)=(-1,-1)\ \text{if }S=\emptyset\quad\text{ and }\quad F(S)=(-x,x)\ \text{if }S=\{x\},$ (11)
Separable embeddings (even non monotone) don’t exist:
Now, as discussed in the main text, we give a justification here on why we choose to relax separability and not monotonicity in the definition of wekaly MAS embeddings. Here, we show that, with the added assumptions of injectivity and continuity, even non-monotone set-t-vector embeddings do not exist in any dimension for uncountable ground sets.
Theorem 16.
There does not exist a continuous injective separable set function taking values in $\mathbb{R}^{m}$ for any $m$ when the ground set $V$ is an open subset of $\mathbb{R}^{d}$
Proof.
We restrict ourselves to $\mathcal{P}_{=n}(V)$, i.e, only subsets of size $n$. Since $F$ is continuous and permutation-invariant, according to Theorem-7 of [20], $f(\{x_{1},\cdots,x_{n}\})=\rho(\sum_{i=1}^{n}\phi(x_{i}))$, where $\rho,\phi$ are continuous and bijective(given in condition). Now, consider the space $\mathcal{S}\subset[0,1]^{n},\mathcal{S}=\{(x_{1},\cdots,x_{n}):x_{1}<x_{2}<x_{3}<\cdots<x_{n};0<x_{i}<1,\forall i\in[n]\}$. Note that $\mathcal{S}$ is open in $0,1]^{n}$ (to see this, let $\epsilon=\min_{i\in[n-1]}(x_{i+1}-x_{i})$. The ball of radius $\frac{\epsilon}{2}$ is contained in the domain). Since $F:\mathcal{S}\to\mathbb{R}^{n}$ is an injective, continuous map. Thus, by Invariance of Domain Theorem([34]), $F(\mathcal{S})$ is open in $\mathbb{R}^{n}$ as well. Note that each element of $S$ uniquely corresponds to an element of $\mathcal{P}_{=n}(V)$. Thus, $F(\mathcal{P}_{=n}(V))$ is open in $\mathbb{R}^{n}$. Now, consider any set $A=\{a_{1},\cdots,a_{n}\}$. Since $F$ is an open map, we have that $\exists\delta>0$ such that $B(F(A),\delta)\subseteq F(\mathcal{P}_{=n}(V))$. Thus, $\exists B\in\mathcal{P}_{=n}(V)$ such that $F(B)_{i}=F(A)_{i}+\frac{\delta}{2n},\forall i\in[n]$. But then, $F(B)>F(A)$ implies $A\subset B$ but $|A|=|B|=n$ by construction. This gives a contradiction. ∎
8.2 Proofs of results on on weakly MAS functions
Proposition 6.
If the deep set model in (2) is implemented with a vector-to-vector monotonously increasing function $\mathrm{M}_{2}$ and a non-negative activation $\sigma$, then $F(\bullet;(A,b))$ is monotone for every $(A,b)$. If in addition $\sigma$ is monotone (increasing or decreasing), then there exists $S\not\subseteq T$ with $F(S;(A,b))\leq F(T;(A,b))$ for all $A,b$.
Proof.
Let $x<y$ and consider the two sets:
$S=\{\frac{x+y}{2}\},T=\{x,y\}$
And we claim that for all $A,b$, although $S\not\subseteq T$, $F(S,(A,b))\leq F(T,(A,b))$. Note that as $x<\frac{x+y}{2}<y$, then it must have been that
$Ax+b\leq A\cdot\frac{x+y}{2}+b\leq Ay+b$
Or
$Ay+b\leq A\cdot\frac{x+y}{2}+b\leq Ax+b$
Then, by (weakly) monotonicity of the activation $\sigma$ and of $M_{2}$, it must be that
$\sigma(A\cdot\frac{x+y}{2}+b)\leq\sigma(A\cdot x+b)$
Or that
$\sigma(A\cdot\frac{x+y}{2}+b)\leq\sigma(A\cdot y+b)$
In any case, it’s true that
$\sigma(A\cdot\frac{x+y}{2})\leq\sigma(A\cdot x+b)+\sigma(A\cdot y+b)$
Then, by monotonicity of $\mathrm{M}_{2}$,
$F(S,(A,b))=\mathrm{M}_{2}\cdot(\sigma(A\cdot\frac{x+y}{2}+b))\leq\mathrm{M}_{2}\cdot(\sigma(A\cdot x+b)+\sigma(A\cdot y+b))=F(T,(A,b))$
Thus,
$F(S,(A,b))\leq F(T,(A,b))$
See 7
Proof.
[…]
Proposition 8.
If $\mathrm{M}_{2}$ is strictly monotone increasing, and $\sigma$ is a hat activation function, then $F(S,(a,b))=\mathrm{M}_{2}(\sum_{x\in S}\sigma(a^{\top}x+b))$ is weakly MAS function.
Proof.
It’s clear the function is a set function. Be $S\subseteq T$, as $\sigma$ is positive, then for any $a,b$,
$\sum_{x\in S}\sigma(a^{T}x+b)\leq\sum_{x\in S}\sigma(a^{T}x+b)+\sum_{x\in T\setminus S}\sigma(a^{T}x+b)=\sum_{x\in T}\sigma(a^{T}x+b)$
So
$F(S,(a,b))\leq F(T,(a,b))$
Denote by $s_{1},s_{2}]$ the support of $\sigma$,and choose some point in it such that $s\in[s_{1},s_{2}]:\sigma(s)>0$. Now, be $S\not\subseteq T$, then there exists $z\in V:C_{S}(z)>C_{T}(z)$. Now, remove $z$ from both sets $C_{T}(z)$ times. This subtracts from both function values the same amount. Now, we have that $z\in S,z\not\in T$. Denote by $\epsilon=\frac{min_{t\in T}|z-t|>0}{2}$. Choose
$a=\frac{s_{2}-s_{1}}{\epsilon},b=-(a\cdot z-s)$
We claim that $\forall t\in T,\sigma(a\cdot t+b)=0$. Assume it’s not zero, thus $s_{1}<a\cdot t+b<s_{2}$. Thus
$s_{1}<\frac{s_{2}-s_{1}}{\epsilon}\cdot t-a\cdot z+s<s_{2}$
Thus,
$s_{1}-s<\frac{s_{2}-s_{1}}{\epsilon}\cdot t-\frac{s_{2}-s_{1}}{\epsilon}\cdot z<s_{2}-s$
Thus,
$\frac{s_{1}-s}{s_{2}-s_{1}}\cdot\epsilon<t-z<\epsilon\cdot\frac{s_{2}-s}{s_{2}-s_{1}}$
But,
$|\frac{s_{1}-s}{s_{2}-s_{1}}|\leq 1$
$|\frac{s_{2}-s}{s_{2}-s_{1}}|\leq 1$
Thus,
$-\epsilon\leq t-z\leq\epsilon$
A contradiction to the definition of $\epsilon$. Thus, $F(T,(a,b))=\mathrm{M}_{2}(0)$. But
$\sigma(a\cdot z+b)=\sigma(a\cdot z-a\cdot z+s_{1})=\sigma(s_{1})>0$
Thus
$F(S,(a,b))=\mathrm{M}_{2}(\sum_{x\in S}\sigma(a^{\top}x+b))\geq\mathrm{M}_{2}(\sigma(s_{1}))>\mathrm{M}_{2}(0)=F(T,(a,b))$
And we are done. ∎
See 10
Proof.
Consider the function $\mathrm{TRI}$ as in the third plot of Fig. 2. We know that $\mathrm{TRI}$ belongs to the class of Hat functions. From 9 we know that, for any $S\not\subseteq T\in\mathcal{P}_{<\infty}(V)$, $\exists\bm{a}\in\mathbb{R}^{d}$ and $b\in\mathbb{R}$ such that $\sum_{x\in S}\mathrm{TRI}(\bm{a}^{\top}x+b)>\sum_{y\in T}\mathrm{TRI}(\bm{a}^{\top}y+b)$. Now, consider the following parameters:
$\begin{split}&\bm{A}=\left[\bm{a},\bm{a},\bm{a}\right]^{\top}\in\mathbb{R}^{3\times d},\bm{b}=[b+1,b-1,b]^{\top}\in\mathbb{R}^{3}\\ &\bm{a}_{1}=[1,1,-2]^{\top}\in\mathbb{R}^{3},\text{ and }b_{1}=0\end{split}$ (12)
Then, for the above values of parameters, we have:
$\mathrm{ReLU}\circ\mathcal{A}_{2}(x)=\left[\mathrm{ReLU}\left((\bm{a}^{\top}x+b)+1\right),\mathrm{ReLU}\left((\bm{a}^{\top}x+b)-1\right),\mathrm{ReLU}\left(\bm{a}^{\top}x+b\right)\right]$
Thus, we have:
$\bm{a}_{1}^{\top}\left(\mathrm{ReLU}\circ\mathcal{A}_{2}(x)\right)+b_{1}$
$=\mathrm{ReLU}\left((\bm{a}^{\top}x+b)+1\right)+\mathrm{ReLU}\left((\bm{a}^{\top}x+b)-1\right)-2\mathrm{ReLU}\left(\bm{a}^{\top}x+b\right)$
$=\mathrm{TRI}\left(\bm{a}^{\top}x+b\right)$
Thus we have, $\mathrm{ReLU}\circ\mathcal{A}_{1}\circ\mathrm{ReLU}\circ\mathcal{A}_{2}(x)=\mathrm{ReLU}\circ\mathrm{TRI}\left(\bm{a}^{\top}x+b\right)=\mathrm{TRI}\left(\bm{a}^{\top}x+b\right)$. Hence, $F(S)=\sum_{x\in S}\mathrm{TRI}\left(\bm{a}^{\top}x+b\right)>\sum_{y\in S}\mathrm{TRI}\left(\bm{a}^{\top}y+b\right)=F(T)$ and weak separability is proved. Since $\mathrm{M}_{2}$ is monotone increasing and $\mathrm{ReLU}$ is non-negative, it follows that $F$ is monotone as well, hence weakly MAS. ∎
8.3 Hölder separability of MAS functions
See 12
Proof.
We consider $V\subset\mathbb{R}^{d}$ to be a compact set with maximum norm of $1$. Also, we note that, if $\sigma(x)$ is a Hat function with support in $\gamma_{1},\gamma_{2}]$, then $\sigma(\frac{x-\gamma_{1}}{\gamma_{2}-\gamma_{1}})$ is a Hat function with support in $(0,1)$. For this proof, we thus consider $\sigma$ to have support in $[0,1]$. For support in general $\gamma_{1},\gamma_{2}]$ and a general norm bound of $V$, the distributions on $b,c$ needs to be scaled and shifted to get same Lower Hölder results with the Hölder constants scaled appropriately. ***If $\sup_{x\in V}\lVert x\rVert\leq B$, and $\text{supp}(\sigma)\subseteq[\gamma_{1},\gamma_{2}]$ then distributions of $b,c$ should be shifted by $\gamma_{1}$ and linearly scaled by $\delta:=\frac{B}{\gamma_{2}-\gamma_{1}}$ and to be: $b\sim\mathrm{Unif}(-\gamma_{1}-\delta,-\gamma_{1}+\delta)$ and $c\sim\mathrm{Unif}(0,2\delta)$ respectively. We recall the definition of $\Delta(S,T)$ from equation 4. Let $S=\left\{x_{1},\cdots,x_{M}\right\}$ and let $T=\left\{y_{1},\cdots,y_{N}\right\}$, where $S,T\subset V$. Clearly by the definition of $\Delta(S,T)$, it is sufficient to consider $|S|\leq|T|$, i.e $M\leq N\leq k$.
Using the definition of EMD gives us the following equivalent definition of $\Delta(S,T)$: If $\Omega_{M,N}:=\left\{\tau:[M]\to[N],\tau\text{ is injective}\right\}$ , then $\Delta(S,T)=\min_{\tau\in\Omega_{M,N}}\sum_{i=1}^{M}\lVert x_{i}-y_{\tau(i)}\rVert_{2}$. Let $\tau^{*}$ be the optimal coupling between $M]\text{ and }[N]$, i.e $\tau^{*}=\arg\min_{\tau\in\Omega_{M,N}}\sum_{i=1}^{M}\lVert x_{i}-y_{\tau(i)}\rVert_{2}$, which means $\Delta(S,T):=\sum_{i=1}^{M}\lVert x_{i}-y_{\tau^{*}(i)}\rVert_{2}$. Now, $\forall i\in[M]$, we define $\Delta_{i}:=\lVert x_{i}-y_{\tau^{*}(i)}\rVert_{2}$. For any $x\in\mathbb{R}^{d},r>0$, let $\mathcal{B}(x,r)$ denote the Euclidean ball centered at $x$ with radius $r$. We define $\Omega_{i}:=\left\{\ell\in[M]:y_{\tau^{*}(\ell)}\in\mathcal{B}(x_{i},\Delta_{i})\right\},\forall i\in[M]$. With these notations, we make the following observation:
Claim-1:
Consider any $r\in(0,k)$. Then, $\exists i_{0}\in[M]$ such that $\sum_{\ell\in\Omega_{i_{0}}}\Delta_{\ell}\leq\Delta_{i_{0}}r$ and $\Delta_{i_{0}}\geq\frac{r^{k}}{k^{k+1}}\Delta(S,T)$
Proof.
For the sake of contradiction, we assume otherwise, i.e $\forall i\in[M]$, we have: either $\Delta_{i}\leq\frac{r^{k}}{k^{k+1}}\Delta(S,T)$ or $\sum_{\ell\in\Omega_{i}}\Delta_{\ell}\geq\Delta_{i}r$. WLOG, assume the following order on $\Delta_{i}$’s: $\Delta_{1}\geq\Delta_{2}\geq\cdots\geq\Delta_{M}$. We build the following directed graph $\mathcal{G}$ on $[M]$. Start with the node $v_{0}=1$ and add we add one node to $\mathcal{G}$ at a time. At step $t$, we select $v_{t}:=\arg\max_{u\in\Omega_{v_{t-1}}}\Delta_{u}$ and add the edge $(v_{t-1},v_{t})$ to $\mathcal{G}$ along with the new node $v_{t}$.
Note that, by definition of $\Delta(S,T)$ we have that: $\Delta(S,T)=\sum_{i=1}^{M}\Delta_{i}$. This, implies that: $\Delta_{v_{0}}\geq\frac{\Delta(S,T)}{M}$, as we take argmax over entire $[M]$ .Thus, $\Delta_{1}\geq\Delta(S,T)/M\geq\Delta(S,T)/k$. Thus, by the contradictory assumption, we must have: $\sum_{u\in\Omega_{v_{0}}}\Delta_{u}\geq\Delta_{1}r$. Since $v_{1}=\arg\max_{u\in\Omega_{v_{0}}}\Delta_{u}$, it thus follows that: $\Delta_{v_{1}}\geq\frac{r}{k}\Delta_{1}\geq\frac{r}{k^{2}}\Delta(S,T)$. Following the similar argument, we thus have, $\Delta_{v_{2}}\geq\frac{r}{k})\Delta_{v_{1}}$ and so on. Thus, $\forall t\in\mathbb{N}$ we must have: $\sum_{u\in\Omega_{v_{t}}}\Delta_{u}\geq\Delta_{v_{t}}r$. Thus, for $v_{t+1}$, defined by: $\arg\max_{u\in\Omega_{v_{t}}}\Delta_{u}$, we must have: $\Delta_{v_{t+1}}\geq\frac{r}{k}\Delta_{v_{t}}$. Since this holds for every $t\in\mathbb{N}$ and since $r>0$ by hypothesis, we must have: $\Delta_{v_{t}}\geq\frac{r}{k}\Delta_{v_{t-1}}\geq\left(\frac{r}{k}\right)^{2}\Delta_{v_{t-1}}\geq\cdots\geq\left(\frac{r}{k}\right)^{t}\Delta_{v_{0}}$.
As we work with finite $M,N$ either $\mathcal{G}$ is a DAG and this process terminates; or $\mathcal{G}$ has a cycle. In former case, at termination step $\bm{t_{t}}$ we cannot add any more nodes, thus $\Omega_{\bm{t_{t}}}=\emptyset$. Since all nodes are unique in this case, we must have $\bm{t_{t}}\leq k$. So, $\Delta_{\bm{t_{t}}}\geq\left(\frac{r}{k}\right)^{\bm{t_{t}}}\Delta_{v_{0}}\geq\frac{1}{k}\left(\frac{r}{k}\right)^{\bm{t_{t}}}\Delta(S,T)$. Since $\frac{r}{k}\in(0,1)$ by hypothesis and since $\bm{t_{t}}\leq k$, we thus have: $\Delta_{\bm{t_{t}}}\geq\frac{1}{k}\left(\frac{r}{k}\right)^{k}\Delta(S,T)$, and hence done with this case.
For the second case, consider the smallest cycle in $\mathcal{G}$. Let this cycle of length $p$ be $\mathcal{C}=\left\{v_{\beta}\to v_{\beta+1}\to\cdots\to v_{\beta+{p-1}}\to v_{\beta}\right\}$ for some $\beta\geq 0$. Then, we have: $\beta+i+1\in\Omega_{\beta+i},\forall i\in\left\{0,\cdots p-1\right\}$, where addition of indices is modulo $p$. Thus we have: $\lVert x_{v_{\beta}+i}-y_{\tau^{*}(v_{\beta}+i+1)}\rVert_{2}<\lVert x_{v_{\beta}+i}-y_{\tau^{*}(v_{\beta}+i)}\rVert_{2},\forall i\in\left\{0,\cdots,p-1\right\}$. Taking sum, we have the following: $\lVert x_{v_{\beta}}-y_{\tau^{*}(v_{\beta}+1)}\rVert_{2}+\lVert x_{v_{\beta}+1}-y_{\tau^{*}(v_{\beta}+2)}\rVert_{2}+\cdots+\lVert x_{v_{\beta}+p-1}-y_{\tau^{*}(v_{\beta})}\rVert_{2}<\sum_{j=0}^{p-1}\lVert x_{v_{\beta}+j}-y_{\tau^{*}(v_{\beta}+j)}\rVert_{2}$. Hence, switching from the coupling $i\mapsto\tau^{*}(i)$ to $i\mapsto\tau^{*}(i+1)$ for $i\in\left\{v_{\beta},\cdots,v_{\beta+p-2}\right\}$ and from $v_{\beta}+p-1\mapsto\tau^{*}(v_{\beta}+p-1)$ to $v_{\beta}+p-1\mapsto\tau^{*}(v_{\beta})$ strictly reduces the value of the sum $\sum_{i=1}^{M}\lVert x_{i}-y_{\tau^{*}(i)}\rVert_{2}$, but we started with the optimal coupling $\tau^{*}$. This gives the reqd. contradiction. ∎
We also make the following observation regarding this "optimal coupling" $\tau^{*}$:
Observation-2:
For any given $i\in[M]$, if $y_{j}\in T\cap\Omega_{i}$ for some $j\in[N]$, then $\exists\ell\in[M]$ such that $j=\tau^{*}(\ell)$
Proof.
The proof follows from the following observation: if $y_{j}\in\Omega_{i}$ for any $i\in[M]$, then we must have: $\lVert x_{i}-y_{j}\rVert_{2}\leq\Delta_{i}=\lVert x_{i}-y_{\tau(i)}\rVert_{2}$. Thus, if $y_{j}$ was "free", i.e if $\not\exists\ell\in[M]:j=\tau^{*}(\ell)$, then we can switch from $i\mapsto\tau^{*}(i)$ to $i\mapsto j$, and this would reduce the sum $\sum_{i=1}^{M}\lVert x_{i}-y_{\tau^{*}(i)}\rVert_{2}$, but $\tau^{*}$ was the optimal coupling. This gives a contradiction ∎
Now, as per the starting discussion, we consider $\text{supp}(\sigma)=[0,1]$, and the proof for general support of $\gamma_{1},\gamma_{2}]$ requires the scaling and shifting as mentioned before.
By assumption we have, $\lim_{t\to 0^{+}}\sigma^{\prime}(t)>0$. We define $\kappa_{1}:=\frac{\lim_{t\to 0^{+}}\sigma^{\prime}(t)}{2}$. Since $\sigma$ is piecewise continuously differentiable as per assumption, we must have $\sigma^{\prime}(t)\geq\kappa_{1},\forall t\in(0,\omega)$ for some $\omega\in(0,1)$ by the continuity of $\sigma^{\prime}$. Note that, by Lagrange’s Mean Value Theorem, we have that:
$\forall x,y\in(0,\omega),\frac{\left|\sigma(x)-\sigma(y)\right|}{\left|x-y\right|}=\sigma^{\prime}(z),\text{ where }z\in(x,y)$
$\implies\forall x,y\in(0,\omega),\left|\sigma(x)-\sigma(y)\right|\geq\kappa_{1}\left|x-y\right|,\text{ since }z\in(0,\omega)$
$\implies\lim_{y\to 0^{+}}\left|\sigma(x)-\sigma(y)\right|\geq\kappa_{1}\left|x\right|,\text{ as }\sigma\text{ is cts.}$
$\implies\sigma(x)\geq\kappa_{1}x,\forall x\in(0,\omega),\text{ as }\sigma\geq 0$
Also, $\sigma$ is upper Lipschitz with constant $\kappa_{2}>0$ per definition of Hat fn. Now, consider the $i_{0}$ coming from the Claim as proved above. We define $\omega^{\prime}:=\min(\frac{\omega}{1-\omega},1)$. Let $P\in\mathbb{N}$ be such that $\frac{\kappa_{1}}{P}<\frac{1}{2}$ and $\frac{\kappa_{1}}{\kappa_{2}P}<k$ and $\omega^{\prime}>\frac{2}{P}$. Note that, such a $P\in\mathbb{N}$ exists, since taking $P\to\infty$ satisfies all the 3 conditions.
Now, for a given $a\in\mathcal{S}^{d-1}$, consider the sets $S_{a}=\left\{a^{\top}x:x\in S\right\},T_{a}=\left\{a^{\top}y:y\in T\right\}$. Note that, since we have $\lVert x_{i}\rVert,\lVert y_{j}\rVert\leq 1,\forall i\in[M],j\in[N]$ as per earlier assumption, we get $S_{a},T_{a}\subseteq(-1,1)$. For given $a\in\mathcal{S}^{d-1}$, we define the "optimal coupling" between the real-valued sets $S_{a},T_{a}$ as: $\tau^{*}_{a}:=\arg\min_{\tau\in\Omega_{M,N}}\sum_{i=1}^{M}\left|a^{\top}x_{i}-a^{\top}y_{\tau(i)}\right|$. Based on this, we define $\Delta^{a}_{i}:=\left|a^{\top}x_{i}-a^{\top}y_{\tau^{*}_{a}(i)}\right|$ and $\Omega^{a}_{i}=\left\{\ell\in[M]:y_{\tau^{*}_{a}(\ell)}\in\mathcal{B}(a^{\top}x_{i},\Delta^{a}_{i})\right\}$. We have: $\Delta(S_{a},T_{a}):=\sum_{i=1}^{M}\Delta^{a}_{i}$. Let $i_{0}\in[M]$ be the index that comes from applying the claim to the sets $S_{a},T_{a}$. Now, we define the regions $\bm{B}_{a}:=(a^{\top}x_{i_{0}}-\omega^{\prime}\Delta^{a}_{i_{0}},a^{\top}x_{i_{0}}-\frac{\Delta^{a}_{i_{0}}}{P})$ and for a given pair $(a\in\mathcal{S}^{d-1},b\in\mathbb{R})$, we consider $\bm{C}_{a,b}:=(\frac{a^{\top}x_{i_{0}}-b}{\omega},a^{\top}x_{i_{0}}+\Delta^{a}_{i_{0}}-b)$. Thus, we have:
$\left(a^{\top}x_{i_{0}}+\Delta^{a}_{i_{0}}-b\right)-\left(\frac{a^{\top}x_{i_{0}}-b}{\omega}\right)$
$=$ $\frac{(1-\omega)(b-a^{\top}x_{i_{0}})+\omega\Delta^{a}_{i_{0}}}{\omega}$
$=$ $\frac{1-\omega}{\omega}\left((b-a^{\top}x_{i_{0}})+\frac{\omega}{1-\omega}\Delta^{a}_{i_{0}}\right)$
$\geq$ $\frac{1-\omega}{\omega}\left(b-(a^{\top}x_{i_{0}}-\omega^{\prime}\Delta^{a}_{i_{0}})\right)>0,\forall b>a^{\top}x_{i_{0}}-\omega^{\prime}\Delta^{a}_{i_{0}}$
The above analysis shows that $\bm{C}_{a,b}\neq\emptyset,\forall b\in\bm{B}_{a}$, so the intervals are well defined. Now, if $b\in\bm{B}_{a}$ and $y<a^{\top}x_{i_{0}}-\Delta^{a}_{i_{0}}$ then we have: $y-b<a^{\top}x_{i_{0}}-\Delta^{a}_{i_{0}}-a^{\top}x_{i_{0}}+\omega^{\prime}\Delta^{a}_{i_{0}}=(\omega^{\prime}-1)\Delta^{a}_{i_{0}}\leq 0$ by definition of $\omega^{\prime}$. On the other hand, if $b\in\bm{B}_{a}$ and $c\in\bm{C}_{a,b}$, for given any $y>a^{\top}x_{i_{0}}+\Delta^{a}_{i_{0}}$ we have: $\frac{y-b}{c}\geq\frac{a^{\top}x_{i_{0}}-b+\Delta^{a}_{i_{0}}}{a^{\top}x_{i_{0}}-b+\Delta^{a}_{i_{0}}}=1$. Combining these, we get that if $\left|y-a^{\top}x_{i_{0}}\right|\geq\Delta^{a}_{i_{0}}$, then $\frac{y-b}{c}\not\in(0,1),\forall b\in\bm{B}_{a},c\in\bm{C}_{a,b}$. Thus, conditioning on a given $a\in\mathcal{S}^{d-1}$, we can write the conditional expectation taken over $b,c$ as follows:
$\mathbb{E}_{b,c}\left[\sum_{x\in S}\sigma\left(\frac{a^{\top}x-b}{c}\right)-\sum_{y\in T}\sigma\left(\frac{a^{\top}y-b}{c}\right)\right]_{+}$
$\geq$ $\mathbb{E}_{b,c}\left[\left(\sum_{x\in S}\sigma\left(\frac{a^{\top}x-b}{c}\right)-\sum_{y\in T}\sigma\left(\frac{a^{\top}y-b}{c}\right)\right)\mathbf{1}_{\{{b\in\bm{B}_{a},c\in\bm{C}_{a,b}}\}}\right]_{+},\text{using non-negative RV}$
$\geq$ $\frac{1}{4}\int_{b=a^{\top}x_{i_{0}}-\omega^{\prime}\Delta^{a}_{i_{0}}}^{a^{\top}x_{i_{0}}}\int_{c=\frac{a^{\top}x_{i_{0}}-b}{\omega}}^{a^{\top}x_{i_{0}}+\Delta^{a}_{i_{0}}-b}\bigg[\sigma\big(\frac{a^{\top}x_{i_{0}}-b}{c}\big)$-
$\sum_{\ell\in\Omega^{a}_{i_{0}}}\bigg|\sigma\big(\frac{a^{\top}y_{\tau^{*}_{a}(\ell)}-b}{c}\big)\ -\sigma\big(\frac{a^{\top}x_{\ell}-b}{c}\big)\bigg|\bigg]_{+}\,\mathrm{d}c\,\mathrm{d}b$
$\geq$ $\frac{1}{4}\int_{b=a^{\top}x_{i_{0}}-\omega^{\prime}\Delta^{a}_{i_{0}}}^{a^{\top}x_{i_{0}}}\int_{c=\frac{a^{\top}x_{i_{0}}-b}{\omega}}^{a^{\top}x_{i_{0}}+\Delta^{a}_{i_{0}}-b}\left[\kappa_{1}\frac{a^{\top}x_{i_{0}}-b}{c}-\kappa_{2}\frac{\kappa_{1}}{P\kappa_{2}}\frac{\Delta^{a}_{i_{0}}}{c}\right]_{+}\,\mathrm{d}c\,\mathrm{d}b,\text{from Lipschitz cts. }\sigma$
[…]
$=$ $\frac{\kappa_{1}}{4}\int_{b=a^{\top}x_{i_{0}}-\omega^{\prime}\Delta^{a}_{i_{0}}}^{a^{\top}x_{i_{0}}-\frac{\Delta^{a}_{i_{0}}}{P}}\left(a^{\top}x_{i_{0}}-b-\frac{\Delta^{a}_{i_{0}}}{P}\right)\ln\left(\frac{\omega(a^{\top}x_{i_{0}}+\Delta^{a}_{i_{0}}-b)}{a^{\top}x_{i_{0}}-b}\right)\,\mathrm{d}b$
$=$ $\frac{\kappa_{1}}{4}\int_{z=\frac{\Delta^{a}_{i_{0}}}{P}}^{\omega^{\prime}\Delta^{a}_{i_{0}}}\left(z-\frac{\Delta^{a}_{i_{0}}}{P}\right)\ln\left(\frac{\omega(z+\Delta^{a}_{i_{0}})}{z}\right)\,\mathrm{d}z,\text{putting }z:=a^{\top}x_{i_{0}}-b$
$\geq$ $\frac{\kappa_{1}}{4}\int_{z=\frac{\Delta^{a}_{i_{0}}}{P}}^{\frac{\omega^{\prime}\Delta^{a}_{i_{0}}}{2}}\left(z-\frac{\Delta^{a}_{i_{0}}}{P}\right)\ln\left(\omega\left(1+\frac{\Delta^{a}_{i_{0}}}{z}\right)\right)\,\mathrm{d}z.\text{using non-negativity of integrand}$
$\geq$ $\frac{\kappa_{1}}{4}\ln\left(\omega\left(1+\frac{2}{\omega^{\prime}}\right)\right)\int_{z=\frac{\Delta^{a}_{i_{0}}}{P}}^{\frac{\omega^{\prime}\Delta^{a}_{i_{0}}}{2}}\left(z-\frac{\Delta^{a}_{i_{0}}}{P}\right)\,\mathrm{d}z$
$\geq$ $\frac{\kappa_{1}}{8}\ln\left(2-\omega\right)\left(\frac{\omega^{\prime}}{2}-\frac{1}{P}\right)^{2}(\Delta^{a}_{i_{0}})^{2},\text{ using }\omega^{\prime}\leq\frac{\omega}{1-\omega}$
[…]
Thus, we have shown that: $\mathbb{E}_{b,c}\bigg[\big[F(S)-F(T)\big]_{+}\big|a\bigg]\geq C^{\prime}\cdot(\Delta^{a}(S,T))^{2}$. By law of total expectation, we have:
$\mathbb{E}_{a,b,c}\bigg[F(S)-F(T)\bigg]_{+}=\mathbb{E}_{a}\mathbb{E}_{b,c}\bigg[\big[F(S)-F(T)\big]_{+}\big|a\bigg]$
$\geq$ $C^{\prime}\mathbb{E}_{a}\big(\Delta^{a}(S,T)\big)^{2}\geq C^{\prime}\left(\mathbb{E}_{a}\Delta^{a}(S,T)\right)^{2},\text{by Jensen's ineq.}$
Thus, to show Lower Hölder separability with exponent $\lambda=2$, it remains to show that, $\mathbb{E}_{a\sim\mathrm{Unif}(\mathcal{S}^{d-1})}\Delta^{a}(S,T)\geq c_{1}\cdot\Delta(S,T)$, where $c_{1}>0$ is a constant. Now, we make the following observation:
Observation-3:
Suppose we’re given given $\ell\in\mathbb{N}$ and $\ell$ many non-zero vectors $u_{1},u_{2},\cdots,u_{\ell}\in\mathbb{R}^{d}$. If $a\sim\mathcal{S}^{d-1}$, then $\exists\delta(\ell)>0$ such that: $\mathbb{P}\left\{\left|a^{\top}u_{i}\right|\geq\delta(\ell)\lVert u_{i}\rVert,\forall i\in[\ell]\right\}\geq\frac{1}{2}$
Proof.
Firstly, we note that, for a fixed $i\in[\ell]$, $\left|a^{\top}u_{i}\right|$ is a continuous non-negative real-valued random variable. Thus, $\exists t_{i}\in\mathbb{R}_{+}$ such that $\mathbb{P}\left\{\left|a^{\top}u_{i}\right|\geq t_{i}\right\}\geq\frac{1}{2}$. Now, Choosing $\delta_{i}:=\frac{t_{i}}{\lVert u_{i}\rVert}>0$ gives us: $\mathbb{P}\left\{\left|a^{\top}u_{i}\right|\geq\delta_{i}\lVert u_{i}\rVert\right\}\geq\frac{1}{2},\forall i\in[\ell]$. Since $\delta_{i}>0,\forall i\in[\ell]$ and $\ell$ is finite, we thus have: $\delta:=\min_{i\in\ell}\delta_{i}>0$. For this particular choice of $\delta$, we have that $\mathbb{P}\left\{\left|a^{\top}u_{i}\right|\geq\delta\lVert u_{i}\rVert\right\}\geq\frac{1}{2},\forall i\in[\ell]$, and the observation is proved. ∎
. Now, consider all $\bm{\ell_{0}}:=\binom{N}{M}M!$ vectors $(x_{i}-y_{\tau(i)})\in\mathbb{R}^{d}$, where $\tau$ runs over all injective functions from $M]\to[N]$. For this particular choice of $\bm{\ell_{0}}$, we get from the last observation that, $\exists\delta(N,M)>0$ such that $\mathbb{P}\left\{\left|a^{\top}\left(x_{i}-y_{\tau(i)}\right)\right|\geq\delta(N,M)\lVert x_{i}-y_{\tau(i)}\rVert\right\}\geq\frac{1}{2}$, for all possible injective $\tau:[M]\to[N]$, we call this set of $a\in\mathcal{S}^{d-1}$ to be $A\subseteq\mathcal{S}^{d-1}$. Now, given a fixed $a\in\mathcal{S}^{d-1}$, we have a particular $\tau_{a}:[M]\to[N]$ that gives $\Delta^{a}(S,T)=\sum_{i=2}^{M}\left|a^{\top}\left(x_{i}-y_{\tau_{a}(i)}\right)\right|$. Also, let $\tau^{*}:[M]\to[N]$ be the "optimal coupling" such that $\Delta(S,T)=\sum_{i=1}^{M}\lVert x_{i}-y_{\tau^{*}(i)}\rVert$. Now, we can write:
$\mathbb{E}_{a\sim\mathrm{Unif}(\mathcal{S}^{d-1})}\Delta^{a}(S,T)$
$=$ $\mathbb{E}_{a\sim\mathrm{Unif}(\mathcal{S}^{d-1})}\left[\sum_{i=1}^{M}\left|a^{\top}(x_{i}-y_{\tau_{a}(i)})\right|\right]$
$\geq$ $\mathbb{E}_{a\sim\mathrm{Unif}(\mathcal{S}^{d-1})}\left[\sum_{i=1}^{M}\left|a^{\top}(x_{i}-y_{\tau_{a}(i)})\right|\bigg|a\in A\right]\mathbb{P}_{a}\left\{a\in A\right\}$
$\geq$ $\frac{\delta(N,M)}{2}\mathbb{E}_{a}\left[\sum_{i=1}^{M}\lVert x_{i}-y_{\tau_{a}(i)}\rVert_{2}\right]$
$\geq$ $\frac{\delta(N,M)}{2}\mathbb{E}_{a}\left[\sum_{i=1}^{M}\lVert x_{i}-y_{\tau^{*}(i)}\rVert_{2}\right],\text{as }\tau^{*}\text{ is optimal coupling}$
$=$ $\frac{\delta(N,M)}{2}\left(\sum_{i=1}^{M}\lVert x_{i}-y_{\tau^{*}(i)}\rVert_{2}\right)=\frac{\delta(N,M)}{2}\Delta(S,T)$
Hence, $F$ is Hölder separable with constant $\frac{C^{\prime}\delta(N,M)}{2}$ and exponent $\lambda=2$. ∎ See 13
Proof.
Denote by $f$ the scalar-valued set function for a co-ordinate of $F$; we proved it’s a lower Holder function. Let $S\not\subseteq T$, then by lower continuity we know that,
$\mathbb{E}_{w\sim\mu(\cdot)}\lVert\left[f(S;w)-f(T;w)\right]_{+}\rVert_{1}\geq C\cdot\Delta(S,T)^{2},\\ \ \text{for all }S,T\in\mathcal{P}_{\leq k}(V)$ (13)
We know that
$\begin{split}&c\cdot\Delta(S,T)^{2}\leq\mathbb{E}_{w\sim\mu(\cdot)}\lVert\left[f(S;w)-f(T;w)\right]_{+}\rVert_{1}\\ \leq&\mathbb{P}(f(S;w)\geq f(T;w))\cdot\sup_{w\in\mathbb{W}}|f(S;w)\geq f(T;w)|\end{split}$
Now, note that as $f$ is a bounded function, denote by $M$ the supremum of $f$ over all sets $S$ and weights $w\in\mathbb{W}$. Thus by the triangle equality,
$c\cdot\Delta(S,T)^{2}\leq 2M\cdot\mathbb{P}(f(S;w)\geq f(T;w))$
Thus, we obtain that
$\frac{c\cdot\Delta(S,T)^{2}}{2\cdot M}\leq\mathbb{P}(f(S;w)\geq f(T;w))$
Thus, the complement satisfies:
$\mathbb{P}(f(S;w)<f(T;w))\leq 1-\frac{c\cdot\Delta(S,T)^{2}}{2\cdot M}$
Now, taking $m$ independeing copies of $f$, namely $F$, we have that
$\mathbb{P}(\forall i\in[m],F(S;w)_{i}<F(T;w)_{i})\leq(1-\frac{c\cdot\Delta(S,T)^{2}}{2\cdot M})^{m}$
Denoting by $C:=\frac{c}{2M}$ yield the desired:
$\mathbb{P}(F(S;w)<F(T;w))\leq(1-C\cdot\Delta(S,T)^{2})^{m}$
Lower Hölder separability of $\mathrm{ReLU}$ networks
So far, we have only given results about Lower Hölder separability of MAS functions with Hat activations. However, as seen in Proposition 10, we have seen that a two layer linear network with $\mathrm{ReLU}$ activation is weakly MAS, and as observed from the proof technique, such a 2 layer $\mathrm{ReLU}$ net can "simulate" a hat activation. In our arguments, we have shown Lower Hölder separability of Hat Activation based parametric set functions by finding a separating parameter point $w\in\mathcal{W}$ for a given pair of sets $(S,T)$, and constructing an open set of non-zero measure, $\mathcal{W}^{\prime}\subseteq\mathcal{W}$ around $w$. We lower bounded the expectation of the non-negative Random variable in the definition of Lower Hölder separability(in  Equation˜5) by restricting the expectation only to $\mathcal{W}^{\prime}$ and showing that it is proportional to $\Delta(S,T)^{\lambda}$. One can extend this exact same idea to construct the open set of parameters for the two later $\mathrm{ReLU}$ networks from the separable parameter point that we get from proof of Proposition 10.
Also, as obtained from Proposition 6, we see that one layer $\mathrm{ReLU}$ networks are not even weakly MAS. But a key key step in that proof uses that one can find 3 collinear elements in the ground set $V$ and construct sets with them to obtain the necessary counter-example. Thus, a natural question is to ask: whether shallow 1-layer networks with $\mathrm{ReLU}$ activation are also weakly MAS with some additional assumptions on the domain. We show in the following result that, with the additional assumption of $V$ being a hypersphere, we may guarantee Lower Hölder separability for even shallow $\mathrm{ReLU}$ networks.
Theorem 17.
Given a ground set $V\subseteq\mathcal{S}^{d-1}$. We define $F:\mathcal{P}_{\leq k}(V)\to\mathbb{R}$ as $F(S)=\sum_{x\in S}\mathrm{ReLU}\left(a^{\top}x+b\right)$, where $a\sim\mathrm{Unif}(\mathcal{S}^{d-1})$ and $b\sim\mathrm{Unif}(-1,1)$. Then, for any $S,T\in\mathcal{P}_{<\infty}(V)$, we have: $\mathbb{E}_{a,b}\left[F(S)-F(T)\right]_{+}\geq C(k)\Delta(S,T)^{(d+3)2^{k}}$ for some constant $C(k)>0$ depending only on $k$.
Proof.
Let $S=\left\{x_{1},\cdots,x_{M}\right\}$ and $T=\left\{y_{1},\cdots,y_{N}\right\}$, where $M\leq N\leq k$. And according to the definition of $\Delta(S,T)$ from Equation˜4 we can equivalently write $\Delta(S,T)=\min_{\pi\in\mathcal{S}_{N}}\sum_{i=1}^{M}\lVert x_{i}-y_{\pi(i)}\rVert_{2}$. Let $\pi^{*}$ be the optimal solution of the above problem, then we can define the map $\tau:[M]\to[N]$ as follows: $\tau(i)=\pi^{*}(i),\forall i\in[M]$. Essentially, $x_{i}$ and $y_{\tau(i)}$ are coupled in the optimal alignment. Now, for any given $i\in[M]$, let $\Delta_{i}:=\lVert x_{i}-y_{\tau(i)}\rVert_{2}$. Then, consider the open Euclidean ball $\mathcal{B}(x_{i},\Delta_{i})$. Then if $j\in[N]$ such that $y_{j}\in\mathcal{B}(x_{i},\Delta_{i})$ then, $j=\tau(\ell)$ for some $\ell\in[M]$. Otherwise switching from $i\mapsto\tau(i)$ to $i\mapsto j$ reduces the cost, which violates optimal alignment. Let $\Omega_{i}:=\left\{\ell\in[M]:y_{\tau(\ell)}\in\mathcal{B}(x_{i},\Delta_{i})\right\}$. Here, we make the following claim:
Claim 18.
$\exists\alpha\in[M]$ such that $\sum_{j\in\Omega_{i}}\lVert y_{j}-x_{\tau^{-1}(j)}\rVert_{2}\leq\frac{\Delta_{\alpha}^{2}}{16}$ and $\Delta_{\alpha}\geq\frac{16\Delta^{2^{k}}(S,T)}{(16k)^{2^{k}}}$
Proof.
Suppose not. WLOG, assume the following order $\Delta_{1}\geq\Delta_{2}\geq\cdots\geq\Delta_{M}$. We build the following directed graph $G$ on $[M]$. Start with the node $v_{0}=1$ and add one node at a time. At step $t$, we select $v_{t}:=\arg\max_{v\in\Omega_{v_{t-1}}}\Delta_{v}$ and add the edge $(v_{t-1},v_{t})$. By the assumption, $\frac{\Delta_{v_{t}}}{16k}\geq(\frac{\Delta_{v_{t-1}}}{16k})^{2}\geq\cdots\geq(\frac{\Delta_{v_{0}}}{(16k)})^{2^{k}}$. Since we work with finite $M$, either $G$ is a DAG and this process terminates or $G$ has a cycle. In former case at termination step $T$ we cannot add any more nodes, thus $\Omega_{T}=\emptyset$, and $T\leq k$. So, $\Delta_{T}\geq\frac{16k}{k}\frac{\Delta^{2^{T}}(S,T)}{(16k)^{2^{T}}}\geq\frac{16\Delta^{2^{k}}(S,T)}{(16k)^{2^{k}}}$ and we’re done. Otherwise, consider the smallest cycle in $G$. Let this cycle of length $p$ be $\mathcal{C}=\left\{v_{\beta}\to v_{\beta+1}\to\cdots\to v_{\beta+{p-1}}\to v_{\beta}\right\}$. Then, we have $\lVert x_{v_{\beta}+i}-y_{\tau(v_{\beta}+i+1)}\rVert_{2}<\lVert x_{v_{\beta}+i}-y_{\tau(v_{\beta}+i)}\rVert_{2},\forall i\in\left\{0,\cdots,p-1\right\}$ and addition of sub-indices is modulo $p$. Thus, we have the following: $\lVert x_{v_{\beta}}-y_{\tau(v_{\beta}+1)}\rVert_{2}+\lVert x_{v_{\beta}+1}-y_{\tau(v_{\beta}+2)}\rVert_{2}+\cdots+\lVert x_{v_{\beta}+p-1}-y_{\tau(v_{\beta})}\rVert_{2}<\sum_{j=0}^{p-1}\lVert x_{v_{\beta}+i}-y_{\tau(v_{\beta}+j)}\rVert_{2}$. Hence, switching from the coupling $i\mapsto\tau(i)$ to $i\mapsto\tau(i+1)$ for $i\in\left\{v_{\beta},\cdots,v_{\beta+p-2}\right\}$ and to $v_{\beta}+p-1\mapsto v_{\beta}$ strictly reduces the cost, but we started with the optimal coupling. This gives a contradiction. ∎
Consider $\alpha$ as in the above claim. Then we define the region: $\mathcal{A}:=\left\{a\in\mathcal{S}^{d-1}:\lVert a-x_{\alpha}\rVert<\frac{\Delta_{\alpha}}{4}\right\}$. Then, $\forall a\in\mathcal{A},\lVert a-x_{\alpha}\rVert_{2}<\lVert a-y_{\tau(\alpha)}\rVert_{2}$. Thus, $a^{\top}x_{\alpha}>a^{\top}y_{\tau(\alpha)}$. Also, consider the region $\mathcal{B}_{a}:=\left\{b\in\mathbb{R}:-1+\frac{\Delta_{\alpha}^{2}}{16}+\frac{\Delta_{\alpha}^{2}}{64}+\frac{\lVert a-x\rVert^{2}}{2}\leq b\leq-1+\frac{\Delta_{\alpha}^{2}}{8}\right\}$. For any given $a\in\mathcal{A}$, we have: $\frac{\lVert a-x\rVert^{2}}{2}\leq\frac{\Delta_{\alpha}^{2}}{32}$, thus $1+\frac{\Delta_{\alpha}^{2}}{16}+\frac{\Delta_{\alpha}^{2}}{64}+\frac{\lVert a-x\rVert^{2}}{2}\leq-1+\frac{7}{64}\Delta_{\alpha}^{2}\leq-1+\frac{\Delta_{\alpha}^{2}}{8}$, so $\mathcal{B}_{a}\neq\emptyset$. Also note that, $\forall a\in\mathcal{A},b\in\mathcal{B}_{a}$, we have:
$\left[F(S)-F(T)\right]_{+}\geq\left[\mathrm{ReLU}\left(a^{\top}x_{\alpha}+b\right)-\left(\sum_{\beta\in\Omega_{\alpha}}\mathrm{ReLU}\left(a^{\top}y_{\tau(\beta)}\right)-\mathrm{ReLU}\left(a^{\top}x_{\beta}\right)\right)\right]_{+}$
$\geq\left[\mathrm{ReLU}\left(a^{\top}x_{\alpha}+b\right)-\sum_{\beta\in\Omega_{\alpha}}\left|\mathrm{ReLU}\left(a^{\top}y_{\tau(\beta)}\right)-\mathrm{ReLU}\left(a^{\top}x_{\beta}\right)\right|\right]_{+}$
$\geq\left[\mathrm{ReLU}\left(a^{\top}x_{\alpha}+b\right)-\sum_{\beta\in\Omega_{\alpha}}\left|a^{\top}(x-y)\right|\right]_{+}$
$\geq\left[\mathrm{ReLU}\left(a^{\top}x_{\alpha}+b\right)-\sum_{\beta\in\Omega_{\alpha}}\lVert x_{\beta}-y_{\tau(\beta)}\rVert\right]_{+}$
$\geq\left[a^{\top}x_{\alpha}+b-\frac{\Delta_{\alpha}^{2}}{16}\right]_{+}=\left[1+b-\frac{\lVert a-x\rVert^{2}}{2}-\frac{\Delta_{\alpha}^{2}}{16}\right]_{+}\geq\frac{\Delta_{\alpha}^{2}}{64}$
Thus, we have:
$\mathbb{E}_{a,b}\left[F(S)-F(T)\right]_{+}\geq\mathbb{E}_{a,b}\left[(F(S)-F(T))\mathbf{1}_{\{{a\in\mathcal{A},b\in\mathcal{B}_{a}}\}}\right]_{+}$
$\geq\frac{\Delta_{\alpha}^{2}}{64}\mathbb{P}\left\{a\in\mathcal{A},b\in\mathcal{B}_{a}\right\}$
$=\frac{\Delta_{\alpha}^{2}}{64S_{d-1}}\int_{a\in\mathcal{S}^{d-1}:a_{1}\geq 1-\frac{\Delta_{\alpha}^{2}}{32}}\frac{1}{2}\left(\frac{3}{64}\Delta_{\alpha}^{2}-(a_{1}-1)\right)da$
$\geq\frac{\Delta_{\alpha}^{4}}{2^{13}S_{d-1}}\int_{a\in\mathcal{S}^{d-1}:a_{1}\geq 1-\frac{\Delta_{\alpha}^{2}}{32}}da$
$=\frac{\Delta_{\alpha}^{4}}{2^{13}}\int_{1-\frac{\Delta_{\alpha}^{2}}{32}}^{1}(1-x^{2})^{\frac{d-3}{2}}dx\geq\frac{\Delta_{\alpha}^{6}}{2^{18}}\left(\frac{\Delta_{\alpha}^{2}}{32}\right)^{\frac{d-3}{2}}\left(2-\frac{\Delta_{\alpha}^{2}}{32}\right)^{\frac{d-3}{2}}$
$\geq\frac{\Delta_{\alpha}^{d+3}}{2^{23}}$
Now, using the bounds from 18, we get that, $\Delta_{\alpha}\geq C(k)\Delta^{2^{k}}(S,T)$. Thus we have that, $\mathbb{E}\left[F(S)-F(T)\right]_{+}\geq C_{1}(k)\Delta(S,T)^{(d+3)2^{k}}$, and done ∎
Upper Lipschitz property of weakly MAS functions
We now show that, the MAS functions obtained by applying one layer neural network with Hat activation for each element of the set, followed by sum aggregation are also Upper Lipschitz continuous in expectation with respect to the augmented Wasserstein distance, following the framework of Davidson and Dym [27]. This shows stability of such MAS embeddings. We say that a parametric set function $F:\mathcal{P}_{\leq k}(V)\times\mathcal{W}\to\mathbb{R}^{m}$ is Upper Lipschitz in expectation if $\exists C>0$ such that:
$\mathbb{E}_{w\in\mathcal{W}}\lVert F(S;w)-F(T;w)\rVert_{1}\leq C\cdot\mathbb{W}^{(k)}(S,T),\ \forall S,T\in\mathcal{P}_{\leq k}(V)$ (14)
Where $\mathbb{W}^{(k)}(.,*)$ is the augmented-Wasserstein metric for sets in $\mathcal{P}_{\leq k}(V)$, where a padding $z$ is added for k|S|k-|S| times to any multiset $S$ of size less than $k$. We work with a compact ground set $V$, and we pick a padding $z$ that has a positive distance from $V$. For the following proofs, we work with $V\subseteq\mathbb{R}^{d}$ which is norm bounded by $1$, and we choose the padding element $z\in\mathbb{R}^{d}$ such that $\lVert z\rVert\geq 3$. For a general $V$ whose norm is bounded by $B$, we can scale the padding $z$ by $B$. With this, we show that, real valued functions $F:\mathcal{P}_{\leq k}(V)\to\mathbb{R}$ given by $F(S)=\sum_{x\in S}\sigma\left(\frac{a^{\top}x-b}{c}\right)$ are Upper Lipschitz w.r.t $\mathbb{W}_{k}$. Now, if we take independent parameteric copies across $m$ output dimensions and if $\mathrm{M}_{2}$ is a vector-to-vector Lipschitz function, then functions of the form $\mathrm{M}_{2}\circ F$, where $F$ has the form in Equation˜6 are Upper Lipschitz as well. This we state in the following theorem:
Theorem 19 ($F$ is upper Lipschitz).
Let $\sigma$ belongs to the class of Hat activations, and $V\subset\mathbb{R}^{d}$ be a compact ground set such that $\sup_{v\in V}\lVert v\rVert\leq 1$. We consider $F:\mathcal{P}_{\leq k}(V)\to\mathbb{R},F(S)=\sum_{x\in S}\sigma\left(\frac{a^{\top}x-b}{c}\right)$, with the distributions $a\sim\mathrm{Unif}(\mathcal{S}^{d-1}),b\sim\mathrm{Unif}(-1,1),c\sim\mathrm{Unif}(0,2)$. Then  $\exists\text{ a constant }C>0$, such that for any $S,T\in\mathcal{P}_{\leq k}(V)$:
$\mathbb{E}_{a,b,c}\left|F(S)-F(T)\right|\leq C\cdot\mathbb{W}^{(k)}(S,T)$
Proof.
Consider $S=\{x_{1},\cdots,x_{M}\},T=\{y_{1},\cdots,y_{N}\}$. Since $\left|F(S)-F(T)\right|$ is symmetric in $S,T$, we consider WLOG that $M\leq N$. As in previous sections, we use $\tau$ to denote an injective function from $M]\to[N]$, and we use $\Omega_{N,M}$ to denote the set of all injective functions from $M]\to[N]$. Thus, we can write the augmented Wasserstein distance on $\mathcal{P}_{\leq k}(V)$ with padding $z$ as:
$\mathbb{W}^{(k)}(S,T)$ $=\arg\min_{\tau\in\Omega_{N,M}}\left\{\sum_{i=1}^{M}\lVert x_{i}-y_{\tau(i)}\rVert+\sum_{j\neq\tau(i)}\lVert y_{j}-z\rVert\right\}$
Let $\tau^{*}$ be the argmin of the above expression that gives the $\mathbb{W}_{k}(S,T)$ Now, we can write:
$\mathbb{E}_{a,b,c}\left|F(S)-F(T)\right|=\mathbb{E}_{a,b,c}\bigg|\sum_{i=1}^{M}\sigma\big(\frac{a^{\top}x_{i}-b}{c}\big)-\sigma\big(\frac{a^{\top}y_{\tau^{*}(i)}-b}{c}\big)-\sum_{j\neq\tau^{*}(i)}\sigma\big(\frac{a^{\top}y_{j}-b}{c}\big)\bigg|$
$\leq$ $\sum_{i=1}^{M}\mathbb{E}_{a,b,c}\left|\sigma\left(\frac{a^{\top}x_{i}-b}{c}\right)-\sigma\left(\frac{a^{\top}y_{\tau^{*}(i)}-b}{c}\right)\right|+\sum_{j\neq\tau^{*}(i)}\mathbb{E}_{a,b,c}\left|\sigma\left(\frac{a^{\top}y_{j}-b}{c}\right)\right|$
We now separately analyze coupled and un-coupled terms: i.e terms of the form $\mathbb{E}\bigg|\sigma\bigg(\frac{a^{\top}y_{j}-b}{c}\bigg)\bigg|$ and $\mathbb{E}\bigg|\sigma(\frac{a^{\top}x_{i}-b}{c})-\sigma(\frac{a^{\top}y_{\tau^{*}(i)}-b}{c})\bigg|$ in the following analysis. For the proof that follows, we consider $\text{supp}(\sigma)\subseteq[0,1]$. For support in $\gamma_{1},\gamma_{2}]$, the proof follows similarly by sifting the distributions of $b,c$ by $\gamma_{1}$ and scaling by $\gamma_{2}-\gamma_{1}$.
Computing $\mathbf{\mathbb{E}\left|\sigma\left(\frac{a^{\top}y_{j}-b}{c}\right)\right|}$:
Note that, $\sigma(\frac{a^{\top}y_{j}-b}{c})=0,\forall b\geq a^{\top}y_{j}$. Also, for a given $b\in(-1,a^{\top}y_{j})$ we have that:
$\forall c\in(0,a^{\top}y_{j}-b),\frac{a^{\top}y_{j}-b}{c}\geq 1\implies\sigma\bigg(\frac{a^{\top}y_{j}-b}{c}\bigg)=0$
Thus, conditioned on a given $a\in\mathcal{S}^{d-1}$, we can write the conditional expectation over $b,c$ as follows:
$\mathbb{E}_{b,c}\left\{\sigma\big(\frac{a^{\top}y_{j}-b}{c}\big)\bigg|a\right\}$
$=$ $\frac{1}{4}\int_{b=-1}^{a^{\top}y_{j}}\int_{c=a^{\top}y_{j}-b}^{2}\bigg|\sigma\bigg(\frac{a^{\top}y_{j}-b}{c}\bigg)-\sigma(0)\bigg|\,\mathrm{d}c\,\mathrm{d}b,\text{ using }\sigma(0)=0$
$\leq$ $\frac{\kappa_{2}}{4}\int_{b=-1}^{a^{\top}y_{j}}\int_{c=a^{\top}y_{j}-b}^{2}\bigg(\frac{a^{\top}y_{j}-b}{c}\bigg)\,\mathrm{d}c\,\mathrm{d}b,\text{as }\sigma\text{ is a {\text{Hat}} activation, it's Lipschitz}$
$=$ $\frac{\kappa_{2}}{4}\int_{b=-1}^{a^{\top}y_{j}}(a^{\top}y_{j}-b)\ln\left(\frac{2}{a^{\top}y_{j}-b}\right)\,\mathrm{d}b=\frac{\kappa_{2}}{4}\int_{a^{\top}y_{j}+1}^{0}z\ln\left(\frac{2}{z}\right)(-\,\mathrm{d}z),z:=a^{\top}y_{j}-b$
$=$ $\frac{\kappa_{2}}{4}\int_{0}^{a^{\top}y_{j}+1}z\ln\left(\frac{2}{z}\right)\,\mathrm{d}z=\frac{\kappa_{2}}{4}\cdot\frac{a^{\top}y_{j}+1}{2}\left(\frac{a^{\top}y_{j}+1}{2}+\left(a^{\top}y_{j}+1\right)\ln\left(\frac{2}{a^{\top}y_{j}+1}\right)\right)$
$\leq$ $\frac{\kappa_{2}}{4}\cdot(a^{\top}y_{j}+a^{\top}a)=\frac{\kappa_{2}}{4}\left|a^{\top}(y_{j}+a)\right|\leq\frac{\kappa_{2}}{4}\lVert y_{j}+a\rVert\leq\frac{\kappa_{2}}{4}\lVert y_{j}-z\rVert$
Now, taking an expectation over $a\sim\mathrm{Unif}(\mathcal{S}^{d-1})$, the inequality is preserved and we get that, for "isolated " $y_{j}$’s, $\mathbb{E}\sigma\left(\frac{a^{\top}-b}{c}\right)\leq\frac{\kappa_{2}}{4}\lVert y_{j}-z\rVert$
Thus for "isolated" elements from the larger set that get coupled with the padded element $z$ while computing augmented-$\mathbb{W}^{(k)}$, we have shown that they are upper bounded by their diatcne from the padding $z$. Hence, it remains to show a similar Lipschitz upper bound for the elements $x_{1},\cdots x_{m}\in S_{1}$ which have a corresponding $y_{i(1)},\cdots,y_{i(m)}\in S_{2}$ in the optimal coupling
Computing $\mathbf{\mathbb{E}\bigg|\sigma(\frac{a^{\top}x_{i}-b}{c})-\sigma(\frac{a^{\top}y_{\tau^{*}(i)}-b}{c})\bigg|}$:
Now, we have, for any $i\in[M]$, and a given $a\in\mathcal{S}^{d-1}$ the conditional expectation $\mathbb{E}_{b,c}\bigg|\sigma(\frac{a^{\top}x_{i}-b}{c})-\sigma(\frac{a^{\top}y_{\tau^{*}(i)}-b}{c})\bigg|$ is symmetric in $x_{i}$ and $y_{\tau^{*}(i)}$. Thus, WLOG we can assume that, $a^{\top}x_{i}\geq a^{\top}y_{\tau^{*}(i)}$. Note that we then have:
$\bigg|\sigma(\frac{a^{\top}x_{i}-b}{c})-\sigma(\frac{a^{\top}y_{\tau^{*}(i)}-b}{c})\bigg|$
$=$ $\begin{cases}&\bigg|\sigma(\frac{a^{\top}x_{i}-b}{c})\bigg|,\forall b\in(a^{\top}y_{\tau^{*}(i)},a^{\top}x_{i}),\forall c\in(a^{\top}x_{i}-b,2)\\ &\bigg|\sigma(\frac{a^{\top}x_{i}-b}{c})-\sigma(\frac{a^{\top}y_{\tau^{*}(i)}-b}{c})\bigg|,\forall b\in(-1,a^{\top}y_{\tau^{*}(i)}),\forall c\in(a^{\top}y_{\tau^{*}(i)}-b,2)\\ &0,\text{ otherwise}\end{cases}$
Hence, conditioned on a specific $a\in\mathcal{S}^{d-1}$, we can write the conditional expectation over $b,c$ as:
$\mathbb{E}_{b,c}\bigg|\sigma(\frac{a^{\top}x_{i}-b}{c})-\sigma(\frac{a^{\top}y_{\tau^{*}(i)}-b}{c})\bigg|=\int_{b=a^{\top}y_{\tau^{*}(i)}}^{a^{\top}x_{i}}\int_{c=a^{\top}x_{i}-b}^{2}\sigma\left(\frac{a^{\top}x_{i}-b}{c}\right)\,\mathrm{d}c\,\mathrm{d}b$
$+\int_{b=-1}^{a^{\top}y_{\tau^{*}(i)}}\int_{c=a^{\top}y_{\tau^{*}(i)}-b}^{2}\bigg|\sigma(\frac{a^{\top}x_{i}-b}{c})-\sigma(\frac{a^{\top}y_{\tau^{*}(i)}-b}{c})\bigg|\,\mathrm{d}c\,\mathrm{d}b$
$\leq$ $\frac{\kappa_{2}}{4}\int_{b=a^{\top}y_{\tau^{*}(i)}}^{a^{\top}x_{i}}(a^{\top}x_{i}-b)\ln\left(\frac{2}{a^{\top}x_{i}-b}\right)\,\mathrm{d}b$
$+\frac{\kappa_{2}}{4}\left|a^{\top}(x_{i}-y_{\tau^{*}(i)})\right|\int_{b=-1}^{a^{\top}y_{\tau^{*}(i)}}\ln\left(\frac{2}{a^{\top}y_{\tau^{*}(i)}-b}\right)\,\mathrm{d}b$
$\leq$ $\frac{\kappa_{2}}{4}\left(\int_{0}^{a^{\top}(x_{i}-y_{\tau^{*}(i)})}z\ln\left(\frac{2}{z}\right)\,\mathrm{d}z+\left|a^{\top}(x_{i}-y_{\tau^{*}(i)})\right|\int_{0}^{a^{\top}y_{\tau^{*}(i)}+1}\ln\left(\frac{2}{z}\right)\,\mathrm{d}z\right)$
$\leq$ $\frac{\kappa_{2}}{4}\left(\left|a^{\top}(x_{i}-y_{\tau^{*}(i)})\right|+2\left|a^{\top}(x_{i}-y_{\tau^{*}(i)})\right|\right)<\kappa_{2}\left|a^{\top}(x_{i}-y_{\tau^{*}(i)})\right|$
$\leq$ $\kappa_{2}\lVert x_{i}-y_{\tau^{*}(i)}\rVert,\text{ by Cauchy Schwarz}$
Thus taking one more expectation over $a\sim\mathcal{S}^{d-1}$, and summing over all $i\in[M]$, we get:
$\mathbb{E}_{a,b,c}\sum_{i=1}^{M}\bigg|\sigma(\frac{a^{\top}x_{i}-b}{c})-\sigma(\frac{a^{\top}y_{\tau^{*}(i)}-b}{c})\bigg|\leq\kappa_{2}\sum_{i=1}^{M}\lVert x_{i}-y_{\tau^{*}(i)}\rVert$
Thus, combining our computations for "isolated" $y_{j}$’s and "coupled" $(x_{i},y_{\tau^{*}(i)})$’s we get that:
$\mathbb{E}_{a,b,c}|F(S)-F(T)|\leq\frac{\kappa_{2}}{4}\sum_{j\neq\tau^{*}(i)}\lVert y_{j}-z\rVert+\kappa_{2}\sum_{i=1}^{M}\lVert x_{i}-y_{\tau^{*}(i)}\rVert\leq\kappa_{2}\mathbb{W}^{(k)}(S,T)$
And our proof of $F$ being upper Lipschitz is complete. ∎
8.4 Universal approximators with MAS functions
See 14
Proof.
A MAS function is in particular invertible, and therefore we can write $f(S)=f\circ F^{-1}\circ F(S)$. On the image of $F$ we can define the function $M=f\circ F^{-1}$ which is a vector-to-vector functions, and it is monotone on the image of $F$, because if $v=F(S),u=F(T)$ for some sets $S,T$, and if $v\leq u$, then by separability $S\subseteq T$, and therefore
$M(v)=f\circ F^{-1}(v)=f(S)\leq f(T)=f\circ F^{-1}(u)=M(u)$
Finally, we need to show that $M$ can be extended to a montone function on all of $\mathbb{R}^{m}$. we accomplish this by defining
$M(v)=\max\{M(u)|\quad u\leq v,u\in\mathrm{Image}(F)\}.$
9 Details about our model
9.1 Architecture details:
In all 4 of MASNet models, we use an elementwise neural network $NN_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{m}$, where the ground set $V\subseteq\mathbb{R}^{d}$ and the output in in $\mathbb{R}^{m}$. We follow that with a sum aggregation, followed by a monotone vector-to-vector neural netwrok $M_{2,\phi}:\mathbb{R}^{m}\to\mathbb{R}^{m}$. For the embedding Neural Network, we use a 1 hidden layer NN with $\mathrm{ReLU}$ activation in the hidden layers and $\mathrm{ReLU}$ or Hat activation in the output, depending on whether we’re using MASNet-ReLU or Hat activation based MASNet. To make $M_{2,\phi}$ monotone, we use non-negative weights by taking the absolute value of parameters before applying the linear transformation, and use all monotonic activations like $\mathrm{ReLU}$ in the intermediate layers. Now, as discussed before, we give details of our model MASNet-INT.
9.2 MASNet-INT
In this appendix we design universal approximators for the class of Hat functions to learn $\sigma$. We parametrize For that, we derive equivalent conditions on the derivatives of $\sigma$ (per embedding dimension) and use the techniques of learning functions by modelling the derivatives using neural networks and then using a numerical integration as in Wehenkel and Louppe [35].
Lemma 20.
If $\sigma:\mathbb{R}\to\mathbb{R}_{\geq 0}$ belongs to the hat activation class , then $\sigma^{\prime}$ satisfies the following conditions for some positive constants $c,C>0$ and $\alpha\in\mathbb{R},\beta>0,\gamma\in(0,1)$:
  1. 1.
    $\text{supp}\left(\sigma^{\prime}\right)\subseteq[\alpha,\alpha+\beta]$
  2. 2.
    $\int_{\alpha}^{\alpha+\beta}\sigma^{\prime}(x)\,\mathrm{d}x=0$
  3. 3.
    $\sigma^{\prime}(x)\leq C,\forall x\in\mathbb{R}$
  4. 4.
    $\sigma^{\prime}(x)\geq c,\forall x\in(\alpha,\alpha+\gamma\cdot\beta)$
The above 4 conditions, along with $\sigma(\alpha)=0$ are the necessary-sufficienet conditions that characterizes the hat activation class.
Proof.
We first prove that, if $\sigma$ belongs to the hat activation class as defined in 8, then conditions 1-4 are satisfied. Note that, if $\sigma$ is a Hat activation, then $\sigma$ has compact support in some $\alpha,\alpha+\beta]$ by deifnition, and it’s piecewise continuously differentiable. Thus, outside $\alpha,\alpha+\beta]$, we have $\sigma\equiv 0\implies\sigma^{\prime}\equiv 0$. Hence, $\text{supp}(\sigma^{\prime})\subseteq[\alpha,\alpha+\beta]$. We also have that, $0=\sigma(\alpha)=\sigma(\alpha+\beta)\implies\sigma(\alpha+\beta)-\sigma(\alpha)=0\implies\int_{\alpha}^{\alpha+\beta}\sigma^{\prime}(t)\,\mathrm{d}t=0$. Finally, $\sigma$ being Lipschitz is equivalent to $\left|\sigma^{\prime}(t)\right|\leq C$ for some $C>0$ and by continuity of $\sigma^{\prime}$ in $(\alpha,\alpha+\beta)$ we must have that $2c:=\lim_{t\to\alpha^{+}}\sigma^{\prime}(t)>0\implies\sigma^{\prime}(t)\geq c,\forall t\in(\alpha,\alpha+\gamma\cdot\beta)$ for some $\gamma\in(0,1)$. Thus, conditions 1-4 are implied by $\sigma$ belonging to the Hat activation class.
Moreover, if conditions 1-4 are satisfied, and $\sigma(\alpha)=0$, then condition-2 implies $\sigma(\alpha+\beta)=0$. This, along with condition-1 implies that $\sigma(t)=0,\forall t\not\in[\alpha,\alpha+\beta]$. Condition-4 immediately implies that $\sigma(\alpha+\gamma\cdot\beta)\geq c\cdot\gamma\beta>0$, thus $\sigma\not\equiv 0$. On the other hand, we get that $\sigma^{\prime}\leq C$ implies by LMVT that, $\left|\sigma(x)-\sigma(y)\right|\leq C\cdot\left|x-y\right|$. Hence, $\sigma$ belongs to the class of Hat activations following the definition from 8. ∎
Neural Parmetrization of MASNet-INT:
Consider $h_{\theta}^{1}(.),h_{\phi}^{2}(.)$ to be non-negative + bounded and bounded fully connected Neural Networks respectively of one hidden layer each. Also consider the trainable support parameters to be $\bm{\alpha}\in\mathbb{R}^{m},\bm{\beta}\in\mathbb{R}^{m}_{+},\bm{\gamma}\in(0,1)^{m}$ where $m$ is the output dimension and $\bm{\alpha},\bm{\beta},\bm{\gamma}$ are the support parameters, aggregated for all output dimensions. Now, we define an integral based model of parametric hat functions. Let $\Theta=(\theta,\phi,\alpha,\beta,m)$ be the parameter space.
$\begin{split}\sigma_{\Theta}(x)&=\int_{\alpha}^{x}h_{\theta}^{1}(z)\mathbf{1}_{\{{\alpha\leq z\leq\alpha+\gamma\beta}\}}\,\mathrm{d}z-\left[\frac{\int_{\alpha}^{\alpha+\gamma\beta}h_{\theta}^{1}(z)dz}{\int_{\alpha+\gamma\beta}^{\alpha+\beta}h_{\phi}^{2}(z)dz}\right]\left(\int_{\alpha+\gamma\beta}^{x}h_{\phi}^{2}(z)\mathbf{1}_{\{{\alpha+\gamma\beta\leq z\leq\alpha+\beta}\}}dz\right)\end{split}$ (15)
In the above formulation, in addition to the support parameters we also learn the function itself rather than performing a linear interpolation, this makes the above class a universal approximator of hat functions.
Lemma 21 (Universal Hat approximator).
The family $\sigma_{\Theta}$ defined in Equation˜15 is an universal approximator of the class of Hat functions having support in $\alpha,\alpha+\beta]$ with $\sigma^{\prime}(t)>0,\forall t\in(\alpha,\alpha+\gamma\beta)$
Proof.
We show that the given parametric model is an universal approximator of Hat functions using their equivalent formulation in Lemma 20. Let $\sigma_{1}$ be $\sigma$ restricted to the interval $\alpha,\alpha+\gamma\beta]$ and $\sigma_{2}$ be $\sigma$ bestricted to the interval $\alpha+\gamma\beta,\alpha+\beta]$. Using Theorem-1 from Lu et al. [36], we can get $\mathrm{ReLU}$ neural network $h_{\theta}^{1}:\mathbb{R}\to\mathbb{R}$ and $h_{\phi}^{2}:\mathbb{R}\to\mathbb{R}$ of width 5 each such that approximates the derivatives of the restricted functions $\sigma_{1}^{\prime}$ and $\sigma_{2}^{\prime}$. Thus, the following hold for any $\epsilon>0$ and any constant $c\in\mathbb{R}$:
$\int_{\alpha}^{\alpha+\gamma\beta}\left|\sigma_{1}^{\prime}(z)-h_{\theta}^{1}(z)\right|\,\mathrm{d}z\leq\epsilon\text{ and }\int_{\alpha+\gamma\beta}^{\alpha+\beta}\left|\sigma_{2}^{\prime}(t)+c\cdot h_{\phi}^{2}(t)\right|\,\mathrm{d}t\leq\epsilon$ (16)
We now observe that, if $x\in(-\infty,\alpha)$, both the indicator functions: $\mathbf{1}_{\{{\alpha\leq z\leq\alpha+\gamma\beta}\}}$ and $\mathbf{1}_{\{{\alpha+\gamma\beta\leq z\leq\alpha+\beta}\}}$ as in the integrands of  Equation˜15 evaluate to 0, thus $\sigma_{\Theta}$ exactly coincides with $\sigma$. On the other hand, if $x\in(\alpha+\beta,\infty)$ then we would have the first integral of 15 evaluate to $\int_{\alpha}^{\alpha+\gamma\beta}h_{\theta}^{1}(z)\,\mathrm{d}z$ (due to presence of the indicator $\mathbf{1}_{\{{\alpha\leq z\leq\alpha+\gamma\beta}\}}$ and the second expression evaluates to $\left[\frac{\int_{\alpha}^{\alpha+\gamma\beta}h_{\theta}^{1}(z)dz}{\int_{\alpha+\gamma\beta}^{\alpha+\beta}h_{\phi}^{2}(z)dz}\right]\left(\int_{\alpha+\gamma\beta}^{\alpha+\beta}h_{\phi}^{2}(z)dz\right)=-\int_{\alpha}^{\alpha+\gamma\beta}h_{\theta}^{1}(z)dz$. Thus, both the expressions sum to 0 and $\sigma_{\Theta}$ coincides with $\sigma$ for $x\in(\alpha+\beta,\infty)$ as well.
Now we shall show that, for any $x\in(\alpha,\alpha+\beta)$ we must also have $|\sigma_{\Theta}(x)-\sigma(x)|\leq\epsilon$, implying convergence in $\sup$ norm. We consider the following cases:
  • Case-I: When $x\in(\alpha,\alpha+\gamma\beta)$, we have that:
    $\left|\sigma(x)-\sigma_{\Theta}(x)\right|$
    $=$ $\left|\int_{\alpha}^{x}\sigma_{1}^{\prime}(z)\,\mathrm{d}z-\int_{\alpha}^{x}h_{\theta}^{1}(z)\,\mathrm{d}z\right|$
    $\leq$ $\int_{\alpha}^{x}\left|\sigma_{1}^{\prime}(z)-h_{\theta}^{1}(z)\right|\,\mathrm{d}z,\text{by traingle ineq.}$
    $\leq$ $\int_{\alpha}^{\alpha+\beta\gamma}\left|\sigma_{1}^{\prime}(z)-h_{\theta}^{1}(z)\right|\,\mathrm{d}z,\text{by non-negativity of integrand}$
[…]
  • Case-II: When $x\in(\alpha+\gamma\beta,\alpha+\beta)$ we have:
    $\left|\sigma(x)-\sigma_{\Theta}(x)\right|$
    $=$ $\left|\int_{\alpha}^{x}\sigma^{\prime}(z)\,\mathrm{d}z-\int_{\alpha}^{\alpha+\gamma\beta}h_{\theta}^{1}(z)\,\mathrm{d}z-\int_{\alpha+\gamma\beta}^{x}h_{\phi}^{2}(z)\,\mathrm{d}z\right|$
    $=$ $\left|\int_{\alpha}^{\alpha+\gamma\beta}\left(\sigma_{1}^{\prime}(z)-h_{\theta}^{1}(z)\right)\,\mathrm{d}z+\int_{\alpha+\gamma\beta}^{x}\left(\sigma_{2}^{\prime}(z)+\left[\frac{\int_{\alpha}^{\alpha+\gamma\beta}h_{\theta}^{1}(z)dz}{\int_{\alpha+\gamma\beta}^{\alpha+\beta}h_{\phi}^{2}(z)dz}\right]h_{\phi}^{2}(z)\right)\,\mathrm{d}z\right|$
    $\leq$ $\int_{\alpha}^{\alpha+\gamma\beta}\left|\sigma_{1}^{\prime}(z)-h_{\theta}^{1}(z)\right|\,\mathrm{d}z+\int_{\alpha+\gamma\beta}^{x}\left|\sigma_{2}^{\prime}(z)+\left[\frac{\int_{\alpha}^{\alpha+\gamma\beta}h_{\theta}^{1}(z)dz}{\int_{\alpha+\gamma\beta}^{\alpha+\beta}h_{\phi}^{2}(z)\,\mathrm{d}z}\right]h_{\phi}^{2}(z)\right|\,\mathrm{d}z,\text{tri. ineq.}$
[…]
    $\leq$ $2\epsilon$
    where last inequality is obtained $c=\frac{\int_{\alpha}^{\alpha+\gamma\beta}h_{\theta}^{1}(z)dz}{\int_{\alpha+\gamma\beta}^{\alpha+\beta}h_{\phi}^{2}(z)\,\mathrm{d}z}$ in the universal approximator for $\sigma_{2}^{\prime}$ in  Equation˜16. Thus, just we see that, for any $\epsilon>0$ and $\alpha\in\mathbb{R},\beta>0,\gamma\in(0,1)$ we can find a universal approximator for the Hat function class
Parametrizing $\alpha,\beta,\gamma$
We have provided an universal approximator for the Hat function class given $\alpha,\beta,\gamma$ but we’d also want to have $\alpha,\beta,\gamma$ as learnable parameters. This is valid for both MASNet-Hat and MASNet-INT, as in both cases we seek to learn the support parameters. Now, there is no constraint on $\alpha$, so we can directly initialize $\bm{\alpha}\in\mathbb{R}^{3m}$ and learn it through gradient descent. However, we have a constraint on $\beta$ that $\beta>0$. For this, we first initialize $\bm{\beta_{0}}\in\mathbb{R}^{m}$ and we obtain $\bm{\beta}\in\mathbb{R}^{m}$ by appling co-ordinatewise positive transformation on $\bm{\beta_{0}}$. This we do in two ways, depending on which works better on a given task: (I) by simply taking $\bm{\beta_{0}}\mapsto\left|\bm{\beta_{0}}\right|$, where the absolute value is taken pointwise. (II)by taking the pointwise transformation $\bm{\beta_{0}}\mapsto\text{ELU}\left({\bm{\beta_{0}}};\upsilon\right)+\upsilon$, where $\text{ELU}\left(\cdot;\upsilon\right)$ is the Exponential Linear Unit with hyperparameter $\upsilon>0$. Note that, the ELU function is given by: $\text{ELU}\left(x;\upsilon\right)=\begin{cases}x,\text{if }x>0\\ \upsilon\left(e^{x}-1\right),\text{if }x\leq 0\end{cases}$. For parametrizing $\gamma\in(0,1)$, we similarly initialize $\bm{\gamma_{0}}\in\mathbb{R}^{m}$ and then apply a pointwise transfprmation that takes each co-ordinate to $(0,1)$. We use tempered sigmoid function: $\bm{\gamma_{0}}\mapsto\text{Sigmoid}\left(\tau\cdot\bm{\gamma_{0}}\right)$ where $\tau>0$ is a hyperparameter
Designing soft indicator functions:
In the integrands for MASNet-INT, we have some indicator functions of the form $\mathbf{1}_{\{{a\leq\cdot\leq b}\}}(x)=\begin{cases}1,\text{if }x\in[a,b]\\ 0,\text{o.w}\end{cases}$. We want these indicator functions to have non zero gradients as we intend to learn the "support parameters" $a,b$ of such indicator function. If we chose binary indicators, then that is not differentiable everywhere and have gradient 0 in most places, making it difficult to learn $a,b$. Thus, we design a "soft" indicator function with the help of tempered sigmoid with hyperparameter $\tau_{S}$ as follows:
$\mathbf{1}_{\{{a\leq\cdot\leq b}\}}(x)=\text{Sigmoid}\left(\tau_{S}\cdot(x-a)\right)\cdot\text{Sigmoid}\left(\tau_{S}\cdot(b-x)\right)$
9.3 Making MASNet Mini Batch Consistent
There has been a chain of works to make set-based neural models to be mini batch consistent in [37, 38] which have applications in large scale machine learning. The goal of these Mini Batch Consistent (MBC) works is to design a set function $F$, which, if applied on any subsets from partitions of a set $S$ and pooled through an activation $g$, should return $F(S)$. Thus, at a high level, MBC functions analyse and design functions for the following batch consistency condition: $g(F(S_{1}),\cdots,F(S_{n}))=F(S)$. MBC methods achieve this by imposing certain restrictions on attention-based set architectures. Similar to this, our work also puts restrictions on the inner embedding transformation and the outer vector-to-vector transformations. It is thus a natural question to ask, if MASNet can be made minibatch consistent. Since set containment is the major theme of our work, with potential applications in retrieval and recommendation systems, processing huge sets at once can be a key bottleneck, and it would be nice to make MASNet process over mini-batches and then design some aggregation mechanism to make the aggregated embedding equal to the embedding of the entire set.
If we consider a partition of a given set $S=\bigcup_{i=1}^{k}S_{i}$, where $S_{i}$’s are all disjoint, then $\textsc{MASNet}(S_{i})\leq\textsc{MASNet}(S)$, by monotonicity. But, we can try to find an appropriate pooling mechanism over the subset embeddings to make the aggregated embedding same as that of the entire set. For example, if the outer transformation $M_{\theta_{2}}$, as in Equation˜7 was invertible, and then we may define a particular permutation-invariant pooling function $g$ such that, $g(X_{1},\cdots,X_{k})=M_{\theta_{2}}\left(\sum_{i=1}^{k}M_{\theta_{2}}^{-1}\left(X_{i}\right)\right)$, then we’d have: $g\left(\textsc{MASNet}(S_{1}),\cdots,\textsc{MASNet}(S_{k})\right)=M_{\theta_{2}}\left(M_{\theta_{2}}^{-1}\left(\textsc{MASNet}(S_{i})\right)\right)=M_{\theta_{2}}\left(\sum_{i=1}^{k}\sum_{x\in S_{i}}\sigma\circ M_{\theta_{1}}(x)\right)=M_{\theta_{2}}\left(\sum_{x\in S}\sigma\circ M_{\theta_{1}}(x)\right)=\textsc{MASNet}(S)$. his is irrespective of how we partition $S$ into $S_{1},\cdots S_{k}$. And this would enable us to process the subsets independently and aggregate them accordingly.
For our purposes, we have needed the outer transformation $M_{\theta_{2}}$ to be a vector-to-vector monotone function in Equation˜7. But that doesn’t guarantee injectivity or invertibility. Thus, we need to restrict the choice of $M_{\theta_{2}}$ to be monotone and invertible for MASNet to be MBC consistent. One possible choice is to use ideas from relevant works like [39, 35] to design monotone and invertible neural networks, and analyse the expressibility properties of the resulting set model. This forms an interesting direction for future work.
10 Additional details about experiments and more experiments
10.1 Hardware details:
All experiments were performed in a compute server, running the OS of GNU Linux Version 12, equipped with a 16 core Intel(R) Xeon(R) Gold 6130 CPU @ 2.10GHz CPU architecture and equipped with a cluster of 6 NVIDIA RTX A6000 GPUs with a memory of 49GB each.
10.2 Dataset details and more experiments
We now give an account of preparation details of each dataset and accounts on more experiments on each.
Synthetic datasets and related experiments:
We begin with a controlled synthetic setting where we generate pairs $(S,T)$ such that $k_{S}<k_{T}$.
Each pair is labeled either positive or negative. For positive examples, we first sample $k_{T}$ vectors in $\mathbb{R}^{d}$ from $\mathcal{N}(0,\mathbb{I}_{d})$, which form the set $T$. A subset of size $k_{S}$ is then uniformly sampled from $T$ to obtain $S$. For negative examples, we again sample $T$ as above, but then independently draw $k_{S}$ vectors from $\mathbb{R}^{d}$ to form $S$ .
To simulate a real world scenerio of information loss, we inject gaussian noise into the vectors of $S$ to get $S^{\prime}$ but keep the boolean label unchanged.
Figure 8: Acc vs $|T|$ for all models $\geq 2$ layer MLP
Figure 9: Acc vs $|S|$ for all models $\geq 2$ layer MLP
On this, we provide two more plots: Figs. 9 and 9 that compares the accuracies of MASNet vs baselines, by varying the gap between target set size $|T|$ and query set size $|S|$. Note that the more this gap is, the more it becomes challenging to separate two non-subsets. In all cases, the embedding MLP $M_{\theta_{1}}$ from Equation˜7 is comprising of $\geq 2$ layers.
Text-based datasets and related experiments:
We evaluate our models on three text based real-world datasets: MSWEB, MSNBC, and the Amazon Baby Registry datasets, all of which reflect naturally occurring set containment structures coming out of user behavior in websites and recommendation engines.
MSWEB and MSNBC:
These datasets consist of user activity logs from www.microsoft.com and www.msnbc.com, respectively. Each user session is treated as a bag of page identifiers, which are embedded into 768-dimensional vectors using a pre-trained BERT model. We construct query-target pairs $(S,T)$ by sampling $S$ from a session $V$, and letting $T=V\setminus S$. A pair is labeled $1$ if $S\subseteq T$, and 0 otherwise. To simulate noisy real-world conditions, Gaussian noise is added to each element in $S$ while preserving the label. The positive-to-negative ratio was taken to be 0.1, which is the number of labels 1 to labels 0 in the entire dataset.
Amazon Baby Registry:
This dataset contains subsets of products selected by customers, each tagged by category (e.g., “toys” or “feeding”). Product descriptions are embedded using BERT. We filter out sets of size <2<2 or >30>30, and for each valid subset $S$, we generate $T\supseteq S$ of size 30 by sampling additional items from the same category. For negative samples, $T$ is sampled randomly and verified to satisfy $S\not\subseteq T$.
Model Bedding Feeding MSWEB MSNBC DeepSets 0.91 0.90 0.93 0.97 SetTransformer 0.91 0.90 0.94 0.96 FlexSubNet 0.90 0.91 0.94 0.94 Neural SFE 0.91 0.91 0.91 0.93 MASNet-ReLU 0.95 0.93 0.97 0.97 MASNet-Hat 0.91 0.92 0.95 0.97
Table 10: Set containment in text datasets with 90:10 ratio
Noise is added to $S$ in both cases. This models scenarios such as predicting whether a set of displayed products $T$ includes a customer’s interest $S$. On these datasets, we provide three sets of additional experiments. In the first table, we perform the same set-containment task as done in the main text, but with a different negative-to-positive ratio of $90:10$.
Note that, due to the inductive bias of monotonicity, all the positive examples are correctly classified by design is case of MASNet. Thus, it only has to learn to identify the to separate the negative examples. Thus, if the test set has a higher proportion of negative examples, then it throws a model a toughter task to learn. This is shown in the following Table 10. Other than the modified class ratio, we provide two additional tables: in the first one, we compare shallow(1 layer) vs deep ($\geq 2$ layers) embedding MLP $M_{\theta_{1}}$ in MASNet, which is shown in Table 11. White noise of std 0.1 added for inexact set containment.
Model bedding feeding MSWEB MSNBC
Shallow-ReLU 0.69 0.36 0.92 0.93
Shallow-MASNet-$\mathrm{TRI}$ 0.83 0.76 0.94 0.95
Shallow-MASNet-Hat 0.89 0.77 0.97 0.96
Shallow-MASNet-INT 0.90 0.91 0.92 0.95
(Deep)MASNet-ReLU 0.95 0.93 0.97 0.97
Deep-MASNet-$\mathrm{TRI}$ 0.92 0.93 0.98 0.96
Deep-MASNet-Hat 0.91 0.92 0.95 0.97
Deep-MASNet-INT 0.93 0.92 0.96 0.97
Table 11: Comparison of shallow vs. deep variants of MASNet across datasets.
Ablation studies on text datasets:
From the definition of MASNet in Equation˜7, we see that, the formulation os quite similar to Deepsets, which is of the form $F(S)=\mathrm{M}_{2}\left(\sum_{x\in S}\mathrm{ReLU}(a^{\top}\mathrm{M}_{1}(x)+b)\right)$. As shown earlier, using $\mathrm{ReLU}$ in the last layer of an elementwise function with a universally approximating $\mathrm{M}_{1}$ is an instance of MASNet.
Model bedding feeding MSWEB MSNBC
DeepSets 0.91 0.90 0.93 0.97
DeepSets, monotone $\mathrm{M}_{2}$ 0.90 0.91 0.94 0.93
MASNet-ReLU 0.95 0.93 0.97 0.97
MASNet-Hat 0.91 0.92 0.95 0.97
MASNet-Hat (No division) 0.87 0.92 0.92 0.94
Table 12: Ablation study
But it is different from DeepSets in a key points, namely: $\bullet$ For set containment tasks, we don’t have an outer $\mathrm{M}_{2}$, which DeepSets have; for universal approximation tasks, we use a monotonially increasing $\mathrm{M}_{2}$ that we enforce by taking positive weights and increasing activation functions. Also, for Hat function based MASNet models, $\bullet$ we use a re-parametrization in which we perform a division based scaling. We now show the effect of the outer monotonic $\mathrm{M}_{2}$ on DeepSets and division based re-parametrization on MASNet-Hat in the ablation study in table 12. White noise of std 0.1 added for inexact containment.
Pointcloud datasets and related experiments
ModelNet40 [33] is a benchmark dataset of 12,311 CAD models across 40 object categories, with each object represented as a 3D point cloud. We frame our task as checking if a given pointcloud $S$ is a segmenet of a target pointcloud $T$.
$|S|\rightarrow$ $128$ $256$ $512$ DeepSets 0.90 0.90 0.91 SetTransformer 0.90 0.91 0.91 FlexSubNet 0.84 0.88 0.92 Neural SFE 0.89 0.89 0.90 MASNet-ReLU 0.91 0.93 0.98 MASNet-Hat 0.87 0.91 0.94
Table 13: Performance on Point cloud for different values of $|S|$, for 90:10 class ratio
Firstly, we choose an object from an object category $C_{1}$, and randomly sample $1024$ points from that object to get the target point cloud $T$. To obtain a positive sample (i.e true subset) from $T$, we first sample a random center point from $T$, and then extract $S$ using a hybrid approach: selecting the nearest point to the center, a few local neighbors via k-NN, and the rest via importance-weighted sampling (inverse-distance from center with noise). This makes sure that we’re selecting a true local region from the point cloud, which is in-line with the actual task of detecting whether a given segment of an object is contained in a target object. For a negative sample(non-subset), we sample $S$ from an object of a different category $C_{2}$.
$|S|\rightarrow$ $128$ $256$ $512$ DeepSets 0.89 0.90 0.90 SetTransformer 0.90 0.91 0.91 MASNet-ReLU 0.87 0.91 0.93 MASNet-Hat 0.85 0.91 0.89 MASNet-$\mathrm{TRI}$ 0.83 0.93 0.87 MASNet-INT 0.87 0.90 0.92
Table 14: Point cloud with DGCNN for different values of $|S|$.
We first give the accuracy table for PointNet encoder, but with the modified negative-to-positive class ratio of 90:10 to make the task harder for MASNet, which correctly classifies the actual subsets by the induictive bias of monotonicity. The numbers with the modified class-ratio are given in the following table of Tab Table˜13:
We also give the accuracy numbers for DGCNN encoder, in Table 14
Details on Linear assignment problem
In this problem, we are given a positive matrix $M\in\mathbb{R}^{n\times m},n\leq m$, where $M_{i,j}$ represents the salary that a ’worker’ will be paid to do a ’job’ $j$. The goal of the task is to maximize the average salary obtained by all workers. This is done by finding the optimal assignment $\pi\in S_{n,m}$ which maps a worker $i\in[n]$ to a ’job’ $\pi(i)\in[m]$, where by construction each worker can be mapped to at most one job, so $\pi(i)\in\{0,1\}$. This gives us the following maximization problem:
$\textstyle F(M)=\frac{1}{n}\max_{\pi\in S_{n,m}}\sum^{n}_{i=1}M_{i,\pi(i)}.$
Thinking of the matrix $M$ as a set of $m$ columns $M=[M_{1},\ldots,M_{m}]$, we see that the function $F$ is permutation invariant and monotone. Accordingly, our goal will be to evaluate our MASNet model and baselines.